### v1

# An Explainable Framework for Particle Swarm Optimization using Landscape Analysis and Machine Learning

Abstract
Swarm intelligence algorithms have demonstrated remarkable success in solving complex optimization problems across diverse domains. However, their widespread adoption is often hindered by limited transparency in how algorithmic components influence performance. This work presents a multi-faceted investigation of Particle Swarm Optimization (PSO) to further understand the key role of different topologies for better interpretability and explainability. To achieve this objective, we first develop a comprehensive landscape characterization framework using Exploratory Landscape Analysis (ELA) to quantify problem difficulty and identify critical features affecting the optimization performance of PSO. Next, we conduct a rigorous empirical study comparing three fundamental swarm communication architectures - Ring, Star, and Von Neumann topologies - analysing their distinct impacts on exploration-exploitation balance, convergence behaviour, and solution quality and eventually develop an explainable benchmarking framework for PSO, to decode how swarm topologies affects information flow, diversity, and convergence. Based on this, a novel machine learning approach for automated algorithm configuration is introduced for training predictive models on extensive Area over the Convergence Curve (AOCC) data to recommend optimal settings based on problem characteristics. Through systematic experimentation across twenty four benchmark functions in multiple dimensions, we establish practical guidelines for topology selection and parameter configuration. These findings advance the development of more transparent and reliable swarm intelligence systems. The source codes of this work can be accessed at https://github.com/GitNitin02/ioh_pso.
keywords:
Particle swarm optimization, Explainable Artificial Intelligence (XAI), Communication Topologies, Hyperparameters analysis, Black box optimization benchmarks
[…]
1 Introduction
Swarm Intelligence (SI) represents a fascinating and powerful field of study that draws inspiration from the collective behaviour observed in decentralized, self-organized natural systems [7]. The complex collective intelligence observed in nature, from the synchronized movements of bird flocks to the optimized foraging of ant colonies, provides the foundational inspiration for sophisticated optimization algorithms. At its core, Swarm Intelligence (SI) is characterized by a population of simple, autonomous agents that follow basic rules and interact locally with their environment and neighbors. Without any centralized guidance, these limited interactions collectively produce sophisticated global patterns and solve complex problems. The approach is fundamentally defined by principles such as decentralization, emergence [7] , adaptation, and positive feedback. Any swarm based optimization algorithm includes three fundamental components: (i) a population that evolves iteratively, (ii) a reproduction process where current population of candidate solutions generates new set of candidate solutions in the next generation, and (iii) a curated information sharing along with an efficient learning mechanism that determines next generation of candidate solutions. These algorithms are inherently scalable, adaptable, and robust, characterized by the simplicity of their individual candidate solutions, making them highly effective for solving a wide array of complex optimization problems [8, 4]. The efficacy of SI algorithms is fundamentally dependent on a critical balance: the exploration of uncharted regions of the search space to discover novel solutions, and the exploitation of known promising areas to refine them. Conversely, exploitation involves refining the search around known promising solutions, leveraging existing information to converge efficiently. Achieving an optimal balance between these two dynamics is paramount for efficiently converging to acceptable solutions while simultaneously avoiding premature convergence to suboptimal local optima. This decentralized architecture is the direct source of the method’s primary strengths: robustness and adaptability [7]. Unlike traditional top-down optimization, the overall solution emerges from the bottom-up interactions of the swarm. This emergent quality makes such systems intrinsically flexible and resilient to dynamic, high-dimensional problem spaces, as the loss of any single agent does not derail the collective process. If one candidate solution fails or a local condition changes, the overall swarm can still adapt and find solutions because its intelligence is distributed, not concentrated in a single vulnerable point. This distributed intelligence makes SI algorithms particularly well-suited for problems where the environment is uncertain, dynamic, or too complex for a single, pre-programmed control mechanism [14]. Among the myriad of SI algorithms, several have gained prominence due to their effectiveness and versatility. Particle Swarm Optimization (PSO) [22] is a prominent algorithm in this category, modeled on the social dynamics of species like birds and fish. In PSO, particles (candidate solutions) traverse a search space, continually adjusting their positions based on their personal best (pbest) known position and the global best (gbest) position discovered by the entire swarm. PSO is particularly well-suited for continuous optimization problems [27], such as the crucial task of parameter tuning [20] in machine learning models. There are several other swarm based optimization algorithms such as artificial bee colony (ABC)[21], ant colony optimization (ACO) [11] and many more but our focus in this article is particle swarm optimization (PSO) which is one of the earliest swarm based optimization algorithms. Parameters play a fundamental role in the effective performance of swarm based optimization algorithms, profoundly influencing a meta-heuristic’s efficiency in solving a given decision problem [2]. The primary function of these parameters is to carefully balance the convergence and diversity of the search process, which directly translates to managing the trade-off between exploration (searching new areas of the solution space) and exploitation (refining the search around currently identified promising solutions) [43]. Suboptimal selection of these control parameters can lead to significant issues, such as premature convergence, where the swarm settles on a suboptimal solution too quickly, and an increase in computational time. The quality of the solution, the amount of computational power needed, the degree of diversity and exploration attained, and the general leadership dynamics within the swarm are all directly impacted by parameter settings [8]. Putting our focus back on PSO, its performance is notably sensitive to its control parameters, particularly the inertia weight $(w)$ and the acceleration coefficients $(c_{1},c_{2})$ [43].Throughout the search process, dynamically balancing exploration and exploitation requires adaptive strategies for these parameters. Our comprehension of PSO parameters is further improved by the following points:
  1. 1.
    Inertia Weight $(w)$: This parameter dictates the momentum of particles, controlling the influence of their previous velocities on their current movement. A higher $w$ encourages particles to explore new regions, promoting global search, while a lower $w$ focuses the search on promising areas, facilitating local search. PSO performance is highly sensitive to the inertia weight.
  2. 2.
    Acceleration Coefficients $(c_{1},c_{2})$: These values are often referred to as “trust parameters,” these coefficients, along with random numbers $(r_{1},r_{2})$, control the stochastic influence of the cognitive (personal best, $c_{1}$) and social (global best, $c_{2}$) components on a particle’s velocity.
    1. (a)
      $c_{1}$ (cognitive component) reflects a particle’s tendency to learn from its own experience and move towards its personal best position.
    2. (b)
      $c_{2}$ (social component) represents a particle’s tendency to follow the collective knowledge of the swarm and move towards the global best position.
    3. (c)
      A proper balance between $c_{1}$ and $c_{2}$ is essential for effective search [10]. If $c_{1}$ is significantly greater than $c_{2}$, particles may wander excessively. Conversely, if $c_{2}$ is significantly greater than $c_{1}$, particles might rush prematurely to local optima.
  3. 3.
    Swarm Size: The total number of particles within the swarm also impacts PSO’s performance.
  4. 4.
    Number of Iterations: While the number of iterations is often just a stopping criterion, studies suggest it has a less substantial impact on performance compared to key parameters such as inertia weight and swarm size [6].
A complex trade-off between convergence speed and the possibility of becoming trapped in local optima is frequently the result of the PSO algorithm’s intrinsic stochasticity and non-linear dynamics, which indicate that parameter settings have a disproportionately large impact on performance [43]. This makes manual tuning a challenging endeavour. Because SI algorithms (including PSO) are stochastic and their search behaviour emerges from simple rules, even minor changes in initial parameter values can result in vastly different search trajectories and outcomes, ranging from divergence to premature convergence or successful optimization. The direct ways to determine optimal parameters PSO are very limited in literature for a given problem, which may have non-linearity and interdependency. This forces researchers to adopt trial-and-error approach, often relying on intuition and extensive experimentation rather than precise calculation [13]. This fundamental challenge of parameter sensitivity is the primary driver for the development of more sophisticated, adaptive parameter strategies, as static and empirically derived parameters are rarely optimal across diverse problem landscapes in the entire optimization process [40].
However, a major challenge lies in the interpretability of these parameters. Their effects are often non-linear and problem-dependent, making it difficult to derive generalizable rules. The next paragraph throws some lights on the need of interpretability and explainability of PSO framework.
1.1 Need of Explainabilty in PSO
Despite its demonstrated efficacy and widespread adoption, PSO, akin to many advanced AI algorithms, it (PSO) frequently functions as a “black box,” [53] obscuring its internal decision-making processes. This inherent lack of clarity poses significant challenges for building trust and verifying the reliability of its outcomes, particularly when applied in sensitive or critical domains like healthcare and autonomous systems. The ambiguous nature of PSO’s configurations and hyperparameters can directly contribute to a perceived lack of reliability in the solutions generated by the algorithm. The stochastic nature of particle movements and the emergent, non-linear interactions within the swarm contribute significantly to this opacity, making it difficult for human users to understand why a specific optimal solution was found or how the swarm navigated the complex search space to converge. This apparent contradiction between PSO’s algorithmic simplicity and its behavioural opacity, reveals a deeper truth: PSO’s algorithmic rules are simple to implement, yet its collective behavior is highly complex and inscrutable. This obscurity does not arise from intricate internal structures but from the swarm’s emergent “intelligence”—the intricate interplay of local and global information sharing across a high-dimensional space. The challenge for XAI is thus that the “black box” nature of such metaheuristics stems from the non-linear dynamics of many simple agents, which confounds a clear tracing of the causal pathway to a solution. Therefore, XAI methods for PSO need to focus on analyzing collective behavior and parameter sensitivity rather than just decomposing complex architectures. In response to this challenge, the fields of Explainable Artificial Intelligence (XAI) and Interpretable Artificial Intelligence (IAI) [12] have rapidly gained prominence. Their primary objective is to develop methodologies that render AI models inherently understandable, thereby clarifying their decision-making processes and nurturing greater confidence in their results.
Historically, metaheuristic optimization algorithms like PSO were primarily evaluated based on their efficiency, convergence speed, and the quality of the solutions they found. However, recent work [16] on analysing the impact of swarm topologies provides an early work in the direction of “enhancing explainability and trustworthiness” in PSO. This work is an extension of the work carried out in [16]. This signals a significant maturation within the field of computational intelligence. It indicates that achieving optimal solutions is no longer the sole criterion for algorithm success. Instead, there is a growing recognition that for AI systems to be responsibly deployed and widely adopted, especially in critical applications, they must also be transparent and understandable [3]. This reflects a broader ethical and practical demand for trustworthy AI. This paradigm shift implies a pressing need for the development of new evaluation metrics that extend beyond traditional optimization performance (e.g., solution quality, computational cost) to quantitatively assess explainability. It also suggests that future research in PSO will increasingly integrate XAI considerations from the very initial design phases of new algorithms, rather than merely applying post-hoc explanations [45].
1.1.1 Analyzing the Impact of Communication Topologies in PSO
The effectiveness and search behavior of PSO are profoundly influenced by its communication topology, which defines the structural pattern through which particles share information within the swarm. Common topologies include Ring, [29, 47] Star [37, 38], and Von Neumann [51, 34]. Research investigates how these distinct topologies impact critical aspects of PSO’s operation, such as information flow, diversity maintenance, convergence speed, and the delicate balance between exploration (searching new regions) and exploitation (refining known good regions). By systematically analyzing these influences, it is possible to enhance the inherent interpretability of PSO’s decision-making processes and derive practical guidelines for selecting the most suitable topology for specific optimization tasks. The explicit detailing of how different communication topologies directly affect PSO’s convergence, diversity, and exploration-exploitation balance provides a powerful form of intrinsic interpretability. This approach is powerful because it moves beyond treating PSO as an indivisible black box. Instead, it focuses on understanding and designing its core components to be inherently transparent. This suggests that modifying the algorithm’s internal architecture and interaction rules can directly contribute to its interpretability, offering a promising avenue for developing more transparent and trustworthy metaheuristics, in future.
1.1.2 Explainable Benchmarking Frameworks for PSO
The development and application of specialized frameworks, such as the adapted IOHxplainer [50], serve as crucial explainable benchmarking tools. These tools can be used to systematically investigate and quantify how different algorithmic configurations, including communication topologies, influence PSO’s performance characteristics. These frameworks are sophisticated enough to support continuous, integer, and categorical parameters, and critically, they account for hyperparameter dependencies, allowing for a detailed analysis of PSO configurations. They often integrate post-hoc XAI methods, such as SHapley Additive exPlanations (SHAP-based techniques) [33], to determine the marginal impact of each parameter on PSO’s performance, thereby significantly aiding in the interpretability of the experimental results.
1.1.3 Understanding Benchmark Problems
Understanding the behaviour of optimization algorithms on benchmark problems is not straightforward, as problem difficulty often depends on hidden structural properties of the search space. Relying only on performance metrics such as convergence speed or solution quality provides limited insight, since these results do not explain why an algorithm succeeds or fails. This is where tools like Exploratory Landscape Analysis (ELA) [35] can play a very crucial role. ELA features allow us to quantify the geometry and structure of optimization problems in a systematic manner. By measuring properties like ruggedness, modality, separability, and variable scaling, ELA helps uncover the reasons behind an algorithm’s performance. For example, highly rugged landscapes may demand more global exploration, while separable problems can be effectively handled by coordinate-wise methods. Such insights are crucial for moving beyond trial-and-error experimentation and towards principled algorithm design and selection. In essence, ELA features have potential to bridge the gap between “black-box” benchmarking and meaningful interpretation. These insights not only guide us in diagnosing problem hardness but also aid in designing adaptive and generalizable optimization strategies.
An metaheuristic optimization algorithm’s performance is largely determined by the particular issue it is trying to solve. To figure out why some problems seem tougher for certain algorithms, we look into how algorithms and problems interact. This approach characterizes optimization problems by extracting meta-features from candidate solutions, which are typically generated using sampling techniques such as Latin hypercube sampling  [44], random sampling, or Sobol sequences [9]. These sampled solutions form the basis for computing landscape features that capture essential problem properties. For explainable analysis in machine learning, several powerful techniques exist to derive these features, these approaches provide a comprehensive framework for understanding and comparing optimization problems in a meaningful way.
Based on the above analysis, the aim of this work is to develop an explainable modelling pipeline that predicts and elucidates the performance of Particle Swarm Optimization (PSO) by linking algorithm hyperparameters to the underlying characteristics of optimization problems. This approach provide deep actionable insights into how specific hyperparameters and their interactions govern PSO’s convergence behaviors, robustness, and adaptability across diverse landscapes. To realize this objective, we design a comprehensive experimental framework that integrates explainable analysis of problem landscapes, systematic exploration of PSO configurations, and interpretable performance modelling. Our experiment is conducted in four integrated parts:
  1. 1.
    We utilize Explainable Landscape Analysis (ELA) to compute and visualize key features (e.g., modality, smoothness, variable interactions) of the Black Box Optimization Benchmark (BBOB) problems, creating an interpretable “fingerprint” for each function’s structure.
  2. 2.
    We construct a comprehensive PSO hyperparameter configuration space containing all necessary hyperparameters for Star, Ring, and Von Neumann topologies.
  3. 3.
    We integrate the configurable PSO into the IOHxplainer framework to generate Area over the Convergence Curve (AOCC) performance data, enabling an analysis of hyperparameter importance and it’s contribution.
  4. 4.
    We train a Decision Tree and Random Forest classifier on the run data to predict performance classes. The resulting tree is visualized to show the feature split at each node, including the distribution of the target class (e.g., high/low performance) for the feature’s values, enabling the identification of critical hyperparameter thresholds and optimal decision rules.
This work is an extension of our own work on explainability of PSO topologies [16]. This paper is organized as follows: Section 2 covers ELA-based visualization for benchmarking. Section 3 proposes a novel explainability framework for PSO. Section 4 presents the experimental setup and results, and Section 5 analyzes the impact of hyperparameters on performance. Section 6 describes our method for data-driven, interpretable configuration learning. Finally, Section 7 concludes and suggests future work.
2 ELA-based visualization for benchmarking
Selecting an effective optimizer requires prior knowledge of a problem’s landscape, as algorithmic performance is heavily influenced by its underlying structure. Traditional high-level characterizations using categorical descriptions of properties like modality, separability, and variable scaling [35] provide a starting point but exhibit critical limitations. They lack quantitative precision, require specialized expertise to interpret, often miss nuanced characteristics, and assume a comprehensive problem understanding that is rarely available in practice.
This gap in understanding is exacerbated by the standard tools of the field. While benchmark suites like the BBOB versions in the COCO platform [18], the CEC function suites [52], and the diverse sets in Nevergrad [42] provide essential platforms for fair and reproducible comparison, they often perpetuate a “black-box” evaluation paradigm. A primary issue is the limited generalizability of results; algorithms that excel on these specific test functions may not perform well on real-world problems. This is frequently due to a lack of benchmark diversity, the use of non-representative baselines, and the fundamental unknown nature of the internal landscape properties that dictate problem difficulty. These limitations can lead to biased comparisons and hinder progress, as the reasons for an algorithm’s success or failure remain opaque.
To overcome the limitations of both high-level classifications and black-box benchmarking, we turn to quantitative, data-driven methods like low-level features [5]. Exploratory Landscape Analysis (ELA) addresses these issues by systematically characterizing problems through sampled data. ELA extracts quantitative meta-features such as convexity, distribution of objective values, local search structure, clustering, and information content from a set of candidate solutions generated via sampling. This provides deeper, more objective insights into the properties that influence algorithm performance.
Tools like the R-package flacco (and its Python counterpart, Pflacco [41]) consolidate this process, offering access to over 300 numerical features across 17 feature sets from a single initial design, a dataset of sampled points and their objective values. The package manages computational overhead by tracking the function evaluations and time required for each feature set. The system delivers essential mathematical function definitions to feature sets that need additional evaluations and supports specialized methods such as cell mapping which needs block divisions across dimensions.
Beyond feature calculation, flacco offers visualization techniques to gain deeper landscape insights. The plotCellMapping [23] function, for instance, visualizes the problem landscape by decomposing the continuous decision space into cells. This method identifies key cell types: attractor cells (black), uncertain cells influenced by multiple attractors (grey), and certain cells that form basins of attraction. Arrows point from each non-attractor cell towards its attractor, with length indicating attraction probability. The interactions and distributions of these cells form the basis for another set of features and provide a powerful visual summary of landscape structure.
Building on these cell mapping foundations, more advanced constructs like barrier trees [25] can be generated. In this representation, local optima (valleys) are depicted as leaves (filled circles), while ridges connecting adjacent valleys are represented by branching nodes (non-filled diamonds), providing a hierarchical model of the fitness landscape’s structure.Visualization of the cell mapping plots and barrier trees plot are shown in Tables 1-3. Similarly, as detailed in Table 4, Information Content of Fitness Sequences (ICoFiS) measures properties such as smoothness, ruggedness, and neutrality by analyzing the entropy in fitness sequences generated from random walks with variable step sizes. This approach effectively adapts discrete information-theoretic measures for continuous optimization domains, with the metrics in Table 4 serving as a direct output of this analysis. The comprehensive nature of this approach is summarized in Tables 5-6, which details feature sets such as those computed by calculate_ela_meta, calculate_ela_distribution, calculate_nbc, calculate_dispersion, and calculate_information_content used to capture the global structure, value distribution, clustering behavior, dispersion, and informational complexity of a selected BBOB function’s landscape. Section 6 details a classification methodology where an expanded suite of ELA features (building on Tables 5 and 6) is used to predict categorical performance of the PSO algorithm, as measured by its AOCC.
Table 1: Analysis of ELA features BBOB functions on d=2.
Function Contour Plot Surface Plot Cell- Mapping Barrier Tree-2D Barrier Tree-3D Information Content
$f_{1}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{2}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{3}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{4}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{5}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{6}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{7}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{8}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 2: Analysis of ELA features BBOB functions on d=2.
Function Contour Plot Surface Plot Cell- Mapping Barrier Tree-2D Barrier Tree-3D Information Content
$f_{9}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{10}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{11}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{12}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{13}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{14}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{15}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{16}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 3: Analysis of ELA features BBOB functions on d=2.
Function Contour Plot Surface Plot Cell- Mapping Barrier Tree-2D Barrier Tree-3D Information Content
$f_{17}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{18}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{19}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{20}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{21}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{22}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{23}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{24}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 4: Information Content visualization of all BBOB functions using iid = 1 on d=5.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{1}$ $f_{2}$ $f_{3}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{4}$ $f_{5}$ $f_{6}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{7}$ $f_{8}$ $f_{9}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{10}$ $f_{11}$ $f_{12}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{13}$ $f_{14}$ $f_{15}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{16}$ $f_{17}$ $f_{18}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{19}$ $f_{20}$ $f_{21}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$f_{22}$ $f_{23}$ $f_{24}$
Table 5: Analysis of ELA features BBOB functions on d=5.
Function iid ela_meta.lin ela_distr ela_distr.num- nbc.nn_nb.cor ic.h_max ic.eps_s ic.eps_max ic.eps_ratio ic.m0
_simple.adj_r2 .skewness ber_of_peaks
$f_{1}$ 1 $6.08\times 10^{-1}$ $5.36\times 10^{-1}$ 2 $6.48\times 10^{-1}$ $8.18\times 10^{-1}$ $9.76\times 10^{-1}$ $1.60\times 10^{-2}$ $1.50\times 10^{0}$ $4.90\times 10^{-1}$
2 $8.60\times 10^{-1}$ $4.01\times 10^{-1}$ 1 $6.65\times 10^{-1}$ $8.18\times 10^{-1}$ $1.08\times 10^{0}$ $1.75\times 10^{-2}$ $1.56\times 10^{0}$ $4.78\times 10^{-1}$
3 $8.32\times 10^{-1}$ $4.22\times 10^{-1}$ 1 $6.48\times 10^{-1}$ $8.10\times 10^{-1}$ $1.06\times 10^{0}$ $1.67\times 10^{-2}$ $1.58\times 10^{0}$ $5.02\times 10^{-1}$
4 $8.09\times 10^{-1}$ $4.12\times 10^{-1}$ 2 $6.54\times 10^{-1}$ $8.12\times 10^{-1}$ $1.10\times 10^{0}$ $1.39\times 10^{-2}$ $1.64\times 10^{0}$ $4.72\times 10^{-1}$
5 $7.91\times 10^{-1}$ $4.60\times 10^{-1}$ 1 $6.47\times 10^{-1}$ $8.10\times 10^{-1}$ $1.08\times 10^{0}$ $1.52\times 10^{-2}$ $1.58\times 10^{0}$ $4.89\times 10^{-1}$
$f_{2}$ 1 $6.63\times 10^{-1}$ $1.03\times 10^{0}$ 2 $5.63\times 10^{-1}$ $7.78\times 10^{-1}$ $7.36\times 10^{-1}$ $1.60\times 10^{-2}$ $1.44\times 10^{0}$ $5.01\times 10^{-1}$
2 $1.80\times 10^{-2}$ $6.45\times 10^{-1}$ 2 $5.40\times 10^{-1}$ $7.84\times 10^{-1}$ $6.36\times 10^{-1}$ $2.31\times 10^{-2}$ $1.22\times 10^{0}$ $5.19\times 10^{-1}$
3 $8.32\times 10^{-1}$ $4.22\times 10^{-1}$ 1 $6.48\times 10^{-1}$ $8.10\times 10^{-1}$ $1.06\times 10^{0}$ $7.29\times 10^{-3}$ $1.74\times 10^{0}$ $5.23\times 10^{-1}$
4 $8.70\times 10^{-1}$ $8.53\times 10^{-1}$ 2 $5.79\times 10^{-1}$ $15\times 10^{-1} 7.7$$76\times 10^{-1}$ $6.96\times 10^{-3}$ $1.66\times 10^{0}$ $5.09\times 10^{-1}$
5 $5.87\times 10^{-3}$ $6.37\times 10^{-1}$ 2 $5.09\times 10^{-1}$ $7.88\times 10^{-1}$ $5.96\times 10^{-1}$ $2.53\times 10^{-2}$ $1.18\times 10^{0}$ $5.15\times 10^{-1}$
$f_{3}$ 1 $5.83\times 10^{-1}$ $6.29\times 10^{-1}$ 1 $6.13\times 10^{-1}$ $8.20\times 10^{-1}$ $8.96\times 10^{-1}$ $1.92\times 10^{-2}$ $1.46\times 10^{0}$ $5.28\times 10^{-1}$
2 $5.87\times 10^{-1}$ $1.96\times 10^{0}$ 2 $6.23\times 10^{-1}$ $7.36\times 10^{-1}$ $5.96\times 10^{-1}$ $3.18\times 10^{-3}$ $1.98\times 10^{0}$ $4.93\times 10^{-1}$
3 $6.97\times 10^{-1}$ $1.44\times 10^{0}$ 2 $6.23\times 10^{-1}$ $6.84\times 10^{-1}$ $6.36\times 10^{-1}$ $2.90\times 10^{-3}$ $1.90\times 10^{0}$ $4.87\times 10^{-1}$
4 $8.55\times 10^{-1}$ $5.49\times 10^{-1}$ 1 $6.35\times 10^{-1}$ $8.00\times 10^{-1}$ $9.56\times 10^{-1}$ $1.83\times 10^{-2}$ $1.58\times 10^{0}$ $4.95\times 10^{-1}$
5 $6.01\times 10^{-1}$ $1.84\times 10^{0}$ 2 $6.21\times 10^{-1}$ $7.24\times 10^{-1}$ $6.56\times 10^{-1}$ $2.77\times 10^{-3}$ $1.94\times 10^{0}$ $5.04\times 10^{-1}$
$f_{4}$ 1 $3.12\times 10^{-1}$ $2.42\times 10^{0}$ 3 $6.10\times 10^{-1}$ $7.55\times 10^{-1}$ $5.16\times 10^{-1}$ $4.01\times 10^{-3}$ $1.96\times 10^{0}$ $5.02\times 10^{-1}$
2 $3.52\times 10^{-1}$ $1.97\times 10^{0}$ 4 $6.37\times 10^{-1}$ $7.39\times 10^{-1}$ $5.96\times 10^{-1}$ $4.19\times 10^{-3}$ $1.84\times 10^{0}$ $5.00\times 10^{-1}$
3 $5.88\times 10^{-1}$ $1.59\times 10^{0}$ 2 $6.13\times 10^{-1}$ $7.23\times 10^{-1}$ $6.56\times 10^{-1}$ $5.28\times 10^{-3}$ $1.70\times 10^{0}$ $4.91\times 10^{-1}$
4 $2.96\times 10^{-1}$ $1.75\times 10^{0}$ 4 $6.16\times 10^{-1}$ $7.82\times 10^{-1}$ $16\times 10^{-1} 6.6$$65\times 10^{-3}$ $1.80\times 10^{0}$ $4.99\times 10^{-1}$
5 $2.63\times 10^{-1}$ $2.93\times 10^{0}$ 4 $6.25\times 10^{-1}$ $7.89\times 10^{-1}$ $5.16\times 10^{-1}$ $2.90\times 10^{-3}$ $2.14\times 10^{0}$ $5.33\times 10^{-1}$
$f_{5}$ 1 $00\times 10^{0} 1.1$$12\times 10^{-1}$ 1 $6.15\times 10^{-1}$ $8.08\times 10^{-1}$ $1.22\times 10^{0}$ $1.21\times 10^{-2}$ $1.66\times 10^{0}$ $4.73\times 10^{-1}$
2 $1.00\times 10^{0}$ $1.68\times 10^{-2}$ 2 $6.14\times 10^{-1}$ $8.03\times 10^{-1}$ $1.20\times 10^{0}$ $1.39\times 10^{-2}$ $1.66\times 10^{0}$ $4.81\times 10^{-1}$
3 $1.00\times 10^{0}$ $4.49\times 10^{-2}$ 2 $6.27\times 10^{-1}$ $8.05\times 10^{-1}$ $1.20\times 10^{0}$ $1.39\times 10^{-2}$ $1.66\times 10^{0}$ $4.79\times 10^{-1}$
4 $1.00\times 10^{0}$ $9.57\times 10^{-2}$ 1 $5.78\times 10^{-1}$ $8.04\times 10^{-1}$ $1.20\times 10^{0}$ $1.46\times 10^{-2}$ $1.66\times 10^{0}$ $4.65\times 10^{-1}$
5 $1.00\times 10^{0}$ $1.04\times 10^{-1}$ 1 $6.14\times 10^{-1}$ $8.05\times 10^{-1}$ $1.20\times 10^{0}$ $1.27\times 10^{-2}$ $1.66\times 10^{0}$ $4.64\times 10^{-1}$
$f_{6}$ 1 $6.13\times 10^{-1}$ $9.78\times 10^{-1}$ 4 $6.42\times 10^{-1}$ $7.82\times 10^{-1}$ $9.36\times 10^{-1}$ $9.62\times 10^{-3}$ $1.70\times 10^{0}$ $4.99\times 10^{-1}$
2 $7.13\times 10^{-1}$ $1.71\times 10^{0}$ 3 $6.01\times 10^{-1}$ $7.26\times 10^{-1}$ $9.56\times 10^{-1}$ $5.53\times 10^{-3}$ $1.86\times 10^{0}$ $4.75\times 10^{-1}$
3 $7.36\times 10^{-1}$ $8.27\times 10^{-1}$ 4 $6.04\times 10^{-1}$ $7.96\times 10^{-1}$ $9.36\times 10^{-1}$ $1.75\times 10^{-2}$ $1.60\times 10^{0}$ $4.77\times 10^{-1}$
4 $8.95\times 10^{-1}$ $7.61\times 10^{-1}$ 2 $6.40\times 10^{-1}$ $7.90\times 10^{-1}$ $1.04\times 10^{0}$ $1.05\times 10^{-2}$ $1.70\times 10^{0}$ $4.69\times 10^{-1}$
5 $9.07\times 10^{-1}$ $7.15\times 10^{-1}$ 3 $6.19\times 10^{-1}$ $7.51\times 10^{-1}$ $9.36\times 10^{-1}$ $1.46\times 10^{-2}$ $1.62\times 10^{0}$ $4.46\times 10^{-1}$
$f_{7}$ 1 $3.80\times 10^{-2}$ $2.03\times 10^{0}$ 4 $6.20\times 10^{-1}$ $8.01\times 10^{-1}$ $9.56\times 10^{-1}$ $1.10\times 10^{-2}$ $1.72\times 10^{0}$ $5.09\times 10^{-1}$
2 $7.80\times 10^{-1}$ $1.01\times 10^{0}$ 1 $6.24\times 10^{-1}$ $7.79\times 10^{-1}$ $1.02\times 10^{0}$ $1.39\times 10^{-2}$ $1.68\times 10^{0}$ $4.79\times 10^{-1}$
3 $7.87\times 10^{-1}$ $1.58\times 10^{0}$ 3 $6.15\times 10^{-1}$ $7.50\times 10^{-1}$ $9.96\times 10^{-1}$ $8.00\times 10^{-3}$ $1.78\times 10^{0}$ $4.74\times 10^{-1}$
4 $3.84\times 10^{-1}$ $1.09\times 10^{0}$ 2 $6.18\times 10^{-1}$ $7.93\times 10^{-1}$ $9.16\times 10^{-1}$ $1.67\times 10^{-2}$ $1.60\times 10^{0}$ $5.33\times 10^{-1}$
5 $2.17\times 10^{-1}$ $1.45\times 10^{0}$ 3 $6.11\times 10^{-1}$ $7.77\times 10^{-1}$ $9.96\times 10^{-1}$ $1.21\times 10^{-2}$ $1.72\times 10^{0}$ $4.88\times 10^{-1}$
$f_{8}$ 1 $3.88\times 10^{-1}$ $9.23\times 10^{-1}$ 2 $6.18\times 10^{-1}$ $7.81\times 10^{-1}$ $7.36\times 10^{-1}$ $1.27\times 10^{-2}$ $1.40\times 10^{0}$ $4.99\times 10^{-1}$
2 $6.55\times 10^{-1}$ $1.20\times 10^{0}$ 1 $6.43\times 10^{-1}$ $7.49\times 10^{-1}$ $7.56\times 10^{-1}$ $1.16\times 10^{-2}$ $1.62\times 10^{0}$ $4.98\times 10^{-1}$
3 $5.99\times 10^{-1}$ $9.97\times 10^{-1}$ 1 $6.34\times 10^{-1}$ $7.57\times 10^{-1}$ $7.96\times 10^{-1}$ $1.52\times 10^{-2}$ $6\times 10^{0} 5.15×1$$5\times 10^{-1}$
4 $5.88\times 10^{-1}$ $1.14\times 10^{0}$ 2 $6.37\times 10^{-1}$ $47\times 10^{-1} 7.7$$76\times 10^{-1}$ $1.60\times 10^{-2}$ $1.54\times 10^{0}$ $4.97\times 10^{-1}$
5 $6.19\times 10^{-1}$ $9.76\times 10^{-1}$ 1 $6.38\times 10^{-1}$ $7.54\times 10^{-1}$ $8.16\times 10^{-1}$ $1.46\times 10^{-2}$ $1.48\times 10^{0}$ $5.05\times 10^{-1}$
$f_{9}$ 1 $5.29\times 10^{-2}$ $2.06\times 10^{0}$ 3 $6.20\times 10^{-1}$ $7.39\times 10^{-1}$ $7.56\times 10^{-1}$ $1.10\times 10^{-2}$ $1.64\times 10^{0}$ $4.69\times 10^{-1}$
2 $6.07\times 10^{-2}$ $1.91\times 10^{0}$ 3 $6.34\times 10^{-1}$ $7.62\times 10^{-1}$ $7.56\times 10^{-1}$ $1.21\times 10^{-2}$ $1.54\times 10^{0}$ $5.30\times 10^{-1}$
3 $5.96\times 10^{-2}$ $2.02\times 10^{0}$ 2 $6.25\times 10^{-1}$ $7.62\times 10^{-1}$ $7.96\times 10^{-1}$ $1.10\times 10^{-2}$ $1.62\times 10^{0}$ $4.97\times 10^{-1}$
4 $7.44\times 10^{-2}$ $3.06\times 10^{0}$ 4 $6.30\times 10^{-1}$ $7.70\times 10^{-1}$ $9.96\times 10^{-1}$ $6.65\times 10^{-3}$ $1.86\times 10^{0}$ $5.17\times 10^{-1}$
5 $5.85\times 10^{-2}$ $2.13\times 10^{0}$ 4 $6.40\times 10^{-1}$ $7.62\times 10^{-1}$ $8.56\times 10^{-1}$ $8.77\times 10^{-3}$ $1.62\times 10^{0}$ $4.97\times 10^{-1}$
$f_{10}$ 1 $6.31\times 10^{-1}$ $1.66\times 10^{0}$ 1 $5.51\times 10^{-1}$ $7.50\times 10^{-1}$ $9.16\times 10^{-1}$ $7.29\times 10^{-3}$ $1.78\times 10^{0}$ $4.95\times 10^{-1}$
2 $6.43\times 10^{-1}$ $1.65\times 10^{0}$ 3 $5.29\times 10^{-1}$ $7.36\times 10^{-1}$ $9.76\times 10^{-1}$ $8.00\times 10^{-3}$ $1.86\times 10^{0}$ $4.69\times 10^{-1}$
3 $7.87\times 10^{-4}$ $1.20\times 10^{0}$ 3 $5.25\times 10^{-1}$ $7.87\times 10^{-1}$ $7.96\times 10^{-1}$ $1.52\times 10^{-2}$ $1.50\times 10^{0}$ $5.17\times 10^{-1}$
4 $7.83\times 10^{-1}$ $1.86\times 10^{0}$ 3 $6.31\times 10^{-1}$ $7.59\times 10^{-1}$ $1.04\times 10^{0}$ $5.04\times 10^{-3}$ $1.96\times 10^{0}$ $4.96\times 10^{-1}$
5 $7.36\times 10^{-1}$ $2.06\times 10^{0}$ 3 $5.42\times 10^{-1}$ $7.28\times 10^{-1}$ $1.02\times 10^{0}$ $5.28\times 10^{-3}$ $1.98\times 10^{0}$ $4.89\times 10^{-1}$
$f_{11}$ 1 $6.20\times 10^{-1}$ $1.60\times 10^{0}$ 2 $5.07\times 10^{-1}$ $7.23\times 10^{-1}$ $8.16\times 10^{-1}$ $7.29\times 10^{-3}$ $1.70\times 10^{0}$ $4.82\times 10^{-1}$
2 $3.63\times 10^{-1}$ $2.11\times 10^{0}$ 3 $5.07\times 10^{-1}$ $7.59\times 10^{-1}$ $9.56\times 10^{-1}$ $8.00\times 10^{-3}$ $1.80\times 10^{0}$ $4.99\times 10^{-1}$
3 $6.81\times 10^{-1}$ $1.65\times 10^{0}$ 3 $5.50\times 10^{-1}$ $7.40\times 10^{-1}$ $9.56\times 10^{-1}$ $9.62\times 10^{-3}$ $1.84\times 10^{0}$ $5.07\times 10^{-1}$
4 $1.90\times 10^{-3}$ $1.73\times 10^{0}$ 4 $4.72\times 10^{-1}$ $7.52\times 10^{-1}$ $8.56\times 10^{-1}$ $1.21\times 10^{-2}$ $1.68\times 10^{0}$ $5.05\times 10^{-1}$
5 $3.20\times 10^{-2}$ $1.52\times 10^{0}$ 3 $4.73\times 10^{-1}$ $7.64\times 10^{-1}$ $8.56\times 10^{-1}$ $1.21\times 10^{-2}$ $1.70\times 10^{0}$ $5.20\times 10^{-1}$
$f_{12}$ 1 $1.06\times 10^{-1}$ $1.94\times 10^{1}$ 4 $6.49\times 10^{-1}$ $7.47\times 10^{-1}$ $1.40\times 10^{0}$ $3.02\times 10^{-5}$ $4.14\times 10^{0}$ $4.93\times 10^{-1}$
2 $4.46\times 10^{-1}$ $5.60\times 10^{0}$ 4 $6.54\times 10^{-1}$ $6.75\times 10^{-1}$ $1.10\times 10^{0}$ $5.78\times 10^{-4}$ $2.62\times 10^{0}$ $4.92\times 10^{-1}$
3 $3.44\times 10^{-1}$ $5.73\times 10^{0}$ 7 $43\times 10^{-1} 6.6$$61\times 10^{-1}$ $8.96\times 10^{-1}$ $1.45\times 10^{-4}$ $2.84\times 10^{0}$ $4.67\times 10^{-1}$
[…]
5 $1.60\times 10^{-1}$ $1.01\times 10^{1}$ 7 $6.44\times 10^{-1}$ $7.35\times 10^{-1}$ $1.04\times 10^{0}$ $1.66\times 10^{-4}$ $3.30\times 10^{0}$ $5.09\times 10^{-1}$
$f_{13}$ 1 $7.44\times 10^{-1}$ $3.28\times 10^{-1}$ 1 $6.26\times 10^{-1}$ $8.03\times 10^{-1}$ $1.10\times 10^{0}$ $1.75\times 10^{-2}$ $1.56\times 10^{0}$ $4.71\times 10^{-1}$
2 $8.34\times 10^{-1}$ $4.79\times 10^{-2}$ 1 $6.49\times 10^{-1}$ $8.18\times 10^{-1}$ $1.16\times 10^{0}$ $1.60\times 10^{-2}$ $1.62\times 10^{0}$ $4.83\times 10^{-1}$
3 $9.60\times 10^{-1}$ $1.97\times 10^{-2}$ 1 $6.63\times 10^{-1}$ $8.03\times 10^{-1}$ $1.16\times 10^{0}$ $1.92\times 10^{-2}$ $1.60\times 10^{0}$ $4.60\times 10^{-1}$
4 $5.53\times 10^{-1}$ $4.65\times 10^{-1}$ 1 $6.24\times 10^{-1}$ $8.09\times 10^{-1}$ $1.06\times 10^{0}$ $1.83\times 10^{-2}$ $1.56\times 10^{0}$ $4.78\times 10^{-1}$
5 $2.67\times 10^{-1}$ $3.61\times 10^{-1}$ 2 $6.30\times 10^{-1}$ $8.20\times 10^{-1}$ $9.96\times 10^{-1}$ $2.01\times 10^{-2}$ $1.52\times 10^{0}$ $4.93\times 10^{-1}$
$f_{14}$ 1 $6.42\times 10^{-1}$ $2.74\times 10^{0}$ 5 $6.40\times 10^{-1}$ $7.08\times 10^{-1}$ $9.76\times 10^{-1}$ $53\times 10^{-3} 2.2$$20\times 10^{0}$ $4.85\times 10^{-1}$
2 $6.47\times 10^{-1}$ $2.52\times 10^{0}$ 3 $6.50\times 10^{-1}$ $7.15\times 10^{-1}$ $9.56\times 10^{-1}$ $1.92\times 10^{-3}$ $2.18\times 10^{0}$ $4.75\times 10^{-1}$
3 $4.60\times 10^{-1}$ $2.58\times 10^{0}$ 4 $6.28\times 10^{-1}$ $7.49\times 10^{-1}$ $9.36\times 10^{-1}$ $4.19\times 10^{-3}$ $1.94\times 10^{0}$ $4.94\times 10^{-1}$
4 $2.98\times 10^{-2}$ $2.89\times 10^{0}$ 4 $6.34\times 10^{-1}$ $7.45\times 10^{-1}$ $8.56\times 10^{-1}$ $3.65\times 10^{-3}$ $1.86\times 10^{0}$ $4.97\times 10^{-1}$
5 $4.20\times 10^{-1}$ $3.32\times 10^{0}$ 4 $6.38\times 10^{-1}$ $7.69\times 10^{-1}$ $9.76\times 10^{-1}$ $4.01\times 10^{-3}$ $2.02\times 10^{0}$ $4.97\times 10^{-1}$
Table 6: Analysis of ELA features BBOB functions on $d=5$.
Function iid ela_meta.lin ela_distr ela_distr.num- nbc.nn_nb.cor ic.h_max ic.eps_s ic.eps_max ic.eps_ratio ic.m0
_simple.adj_r2 .skewness ber_of_peaks
$f_{15}$ 1 $5.56\times 10^{-1}$ $2.38\times 10^{0}$ 3 $6.44\times 10^{-1}$ $7.52\times 10^{-1}$ $8.36\times 10^{-1}$ $3.33\times 10^{-3}$ $2.02\times 10^{0}$ $4.89\times 10^{-1}$
2 $5.04\times 10^{-1}$ $1.40\times 10^{0}$ 1 $6.07\times 10^{-1}$ $8.13\times 10^{-1}$ $9.56\times 10^{-1}$ $1.16\times 10^{-2}$ $1.70\times 10^{0}$ $4.93\times 10^{-1}$
3 $3.70\times 10^{-1}$ $2.05\times 10^{0}$ 5 $6.09\times 10^{-1}$ $8.21\times 10^{-1}$ $8.96\times 10^{-1}$ $7.29\times 10^{-3}$ $1.84\times 10^{0}$ $5.22\times 10^{-1}$
4 $4.28\times 10^{-1}$ $3.56\times 10^{0}$ 5 $6.10\times 10^{-1}$ $7.67\times 10^{-1}$ $8.16\times 10^{-1}$ $3.49\times 10^{-3}$ $2.10\times 10^{0}$ $5.06\times 10^{-1}$
5 $3.16\times 10^{-1}$ $2.14\times 10^{0}$ 3 $5.83\times 10^{-1}$ $7.92\times 10^{-1}$ $1.10\times 10^{0}$ $6.06\times 10^{-3}$ $1.90\times 10^{0}$ $5.26\times 10^{-1}$
$f_{16}$ 1 $3.72\times 10^{-3}$ $1.12\times 10^{0}$ 2 $3.37\times 10^{-1}$ $8.75\times 10^{-1}$ $4.35\times 10^{-1}$ $3.50\times 10^{-2}$ $1.18\times 10^{0}$ $6.51\times 10^{-1}$
2 $5.17\times 10^{-3}$ $9.31\times 10^{-1}$ 1 $3.96\times 10^{-1}$ $8.57\times 10^{-1}$ $4.75\times 10^{-1}$ $4.01\times 10^{-2}$ $1.12\times 10^{0}$ $6.39\times 10^{-1}$
3 $5.42\times 10^{-4}$ $1.01\times 10^{0}$ 2 $4.04\times 10^{-1}$ $8.58\times 10^{-1}$ $4.35\times 10^{-1}$ $3.50\times 10^{-2}$ $1.18\times 10^{0}$ $6.68\times 10^{-1}$
4 $4.41\times 10^{-4}$ $8.85\times 10^{-1}$ 2 $3.91\times 10^{-1}$ $8.70\times 10^{-1}$ $4.15\times 10^{-1}$ $3.83\times 10^{-2}$ $1.14\times 10^{0}$ $6.78\times 10^{-1}$
5 $3.38\times 10^{-3}$ $1.15\times 10^{0}$ 2 $3.58\times 10^{-1}$ $8.60\times 10^{-1}$ $4.35\times 10^{-1}$ $3.19\times 10^{-2}$ $1.20\times 10^{0}$ $6.59\times 10^{-1}$
$f_{17}$ 1 $5.31\times 10^{-1}$ $2.78\times 10^{0}$ 4 $5.24\times 10^{-1}$ $7.70\times 10^{-1}$ $6.96\times 10^{-1}$ $3.33\times 10^{-3}$ $2.14\times 10^{0}$ $5.51\times 10^{-1}$
2 $2.69\times 10^{-1}$ $4.64\times 10^{0}$ 5 $5.11\times 10^{-1}$ $8.34\times 10^{-1}$ $7.96\times 10^{-1}$ $5.53\times 10^{-3}$ $2.04\times 10^{0}$ $5.95\times 10^{-1}$
3 $3.17\times 10^{-1}$ $4.90\times 10^{0}$ 7 $5.07\times 10^{-1}$ $8.30\times 10^{-1}$ $9.16\times 10^{-1}$ $3.49\times 10^{-3}$ $2.12\times 10^{0}$ $5.85\times 10^{-1}$
4 $5.13\times 10^{-1}$ $2.33\times 10^{0}$ 5 $5.43\times 10^{-1}$ $7.74\times 10^{-1}$ $7.56\times 10^{-1}$ $6.96\times 10^{-3}$ $1.92\times 10^{0}$ $5.65\times 10^{-1}$
5 $3.30\times 10^{-1}$ $2.12\times 10^{0}$ 4 $4.94\times 10^{-1}$ $8.64\times 10^{-1}$ $7.56\times 10^{-1}$ $1.60\times 10^{-2}$ $1.56\times 10^{0}$ $6.22\times 10^{-1}$
$f_{18}$ 1 $4.70\times 10^{-1}$ $3.59\times 10^{0}$ 4 $5.31\times 10^{-1}$ $7.97\times 10^{-1}$ $7.56\times 10^{-1}$ $3.04\times 10^{-3}$ $2.20\times 10^{0}$ $6.03\times 10^{-1}$
2 $2.80\times 10^{-1}$ $3.62\times 10^{0}$ 6 $5.24\times 10^{-1}$ $8.31\times 10^{-1}$ $8.16\times 10^{-1}$ $5.53\times 10^{-3}$ $2.00\times 10^{0}$ $5.99\times 10^{-1}$
3 $3.43\times 10^{-1}$ $4.68\times 10^{0}$ 6 $4.92\times 10^{-1}$ $8.13\times 10^{-1}$ $8.16\times 10^{-1}$ $3.65\times 10^{-3}$ $2.12\times 10^{0}$ $6.19\times 10^{-1}$
4 $4.10\times 10^{-1}$ $3.73\times 10^{0}$ 5 $5.52\times 10^{-1}$ $7.81\times 10^{-1}$ $9.36\times 10^{-1}$ $3.65\times 10^{-3}$ $2.12\times 10^{0}$ $5.65\times 10^{-1}$
5 $3.39\times 10^{-1}$ $1.47\times 10^{0}$ 3 $5.13\times 10^{-1}$ $8.56\times 10^{-1}$ $6.96\times 10^{-1}$ $2.11\times 10^{-2}$ $1.44\times 10^{0}$ $6.22\times 10^{-1}$
$f_{19}$ 1 $7.35\times 10^{-2}$ $2.02\times 10^{0}$ 3 $5.67\times 10^{-1}$ $7.97\times 10^{-1}$ $8.16\times 10^{-1}$ $1.16\times 10^{-2}$ $1.66\times 10^{0}$ $5.62\times 10^{-1}$
2 $5.20\times 10^{-2}$ $2.12\times 10^{0}$ 3 $5.61\times 10^{-1}$ $8.01\times 10^{-1}$ $7.76\times 10^{-1}$ $1.05\times 10^{-2}$ $1.64\times 10^{0}$ $5.09\times 10^{-1}$
3 $5.09\times 10^{-2}$ $1.70\times 10^{0}$ 4 $5.50\times 10^{-1}$ $7.93\times 10^{-1}$ $7.96\times 10^{-1}$ $1.21\times 10^{-2}$ $1.58\times 10^{0}$ $5.28\times 10^{-1}$
4 $7.16\times 10^{-2}$ $1.93\times 10^{0}$ 3 $5.56\times 10^{-1}$ $98\times 10^{-1} 7.7$$76\times 10^{-1}$ $1.01\times 10^{-2}$ $1.72\times 10^{0}$ $5.61\times 10^{-1}$
5 $2.26\times 10^{-2}$ $3.04\times 10^{0}$ 5 $5.03\times 10^{-1}$ $7.96\times 10^{-1}$ $8.96\times 10^{-1}$ $8.00\times 10^{-3}$ $1.74\times 10^{0}$ $5.23\times 10^{-1}$
$f_{20}$ 1 $6.56\times 10^{-1}$ $1.40\times 10^{0}$ 1 $6.29\times 10^{-1}$ $7.74\times 10^{-1}$ $8.76\times 10^{-1}$ $1.10\times 10^{-2}$ $1.66\times 10^{0}$ $5.17\times 10^{-1}$
2 $6.69\times 10^{-1}$ $1.42\times 10^{0}$ 2 $6.24\times 10^{-1}$ $7.78\times 10^{-1}$ $8.96\times 10^{-1}$ $1.16\times 10^{-2}$ $1.72\times 10^{0}$ $4.97\times 10^{-1}$
3 $6.49\times 10^{-1}$ $1.47\times 10^{0}$ 2 $6.22\times 10^{-1}$ $7.67\times 10^{-1}$ $9.16\times 10^{-1}$ $9.18\times 10^{-3}$ $1.70\times 10^{0}$ $4.99\times 10^{-1}$
4 $6.68\times 10^{-1}$ $1.36\times 10^{0}$ 1 $6.06\times 10^{-1}$ $7.79\times 10^{-1}$ $8.76\times 10^{-1}$ $1.16\times 10^{-2}$ $1.68\times 10^{0}$ $5.01\times 10^{-1}$
5 $6.57\times 10^{-1}$ $1.50\times 10^{0}$ 2 $6.25\times 10^{-1}$ $7.79\times 10^{-1}$ $8.96\times 10^{-1}$ $9.18\times 10^{-3}$ $1.72\times 10^{0}$ $5.30\times 10^{-1}$
$f_{21}$ 1 $1.51\times 10^{-2}$ $1.28\times 10^{0}$ 2 $4.13\times 10^{-1}$ $8.21\times 10^{-1}$ $5.16\times 10^{-1}$ $2.65\times 10^{-2}$ $1.28\times 10^{0}$ $5.95\times 10^{-1}$
2 $25\times 10^{-2} 1.1$$15\times 10^{0}$ 1 $4.39\times 10^{-1}$ $8.25\times 10^{-1}$ $4.95\times 10^{-1}$ $2.42\times 10^{-2}$ $1.26\times 10^{0}$ $5.90\times 10^{-1}$
3 $2.67\times 10^{-2}$ $1.16\times 10^{0}$ 1 $4.15\times 10^{-1}$ $8.26\times 10^{-1}$ $4.75\times 10^{-1}$ $2.91\times 10^{-2}$ $1.20\times 10^{0}$ $5.89\times 10^{-1}$
4 $9.02\times 10^{-3}$ $1.25\times 10^{0}$ 1 $3.69\times 10^{-1}$ $8.06\times 10^{-1}$ $5.16\times 10^{-1}$ $2.91\times 10^{-2}$ $1.22\times 10^{0}$ $5.70\times 10^{-1}$
5 $4.01\times 10^{-2}$ $1.12\times 10^{0}$ 2 $4.93\times 10^{-1}$ $8.11\times 10^{-1}$ $4.95\times 10^{-1}$ $2.31\times 10^{-2}$ $1.22\times 10^{0}$ $6.03\times 10^{-1}$
$f_{22}$ 1 $6.65\times 10^{-2}$ $2.71\times 10^{0}$ 5 $4.70\times 10^{-1}$ $7.92\times 10^{-1}$ $4.75\times 10^{-1}$ $1.10\times 10^{-2}$ $1.54\times 10^{0}$ $5.83\times 10^{-1}$
2 $4.71\times 10^{-2}$ $2.23\times 10^{0}$ 3 $4.84\times 10^{-1}$ $7.94\times 10^{-1}$ $4.95\times 10^{-1}$ $1.33\times 10^{-2}$ $1.52\times 10^{0}$ $5.66\times 10^{-1}$
3 $5.01\times 10^{-2}$ $2.20\times 10^{0}$ 3 $4.58\times 10^{-1}$ $7.93\times 10^{-1}$ $4.95\times 10^{-1}$ $1.46\times 10^{-2}$ $1.52\times 10^{0}$ $5.50\times 10^{-1}$
4 $4.30\times 10^{-2}$ $2.06\times 10^{0}$ 2 $4.59\times 10^{-1}$ $7.81\times 10^{-1}$ $4.55\times 10^{-1}$ $1.27\times 10^{-2}$ $1.54\times 10^{0}$ $5.53\times 10^{-1}$
5 $1.07\times 10^{-1}$ $2.34\times 10^{0}$ 3 $5.12\times 10^{-1}$ $7.85\times 10^{-1}$ $5.56\times 10^{-1}$ $1.16\times 10^{-2}$ $1.58\times 10^{0}$ $5.64\times 10^{-1}$
$f_{23}$ 1 $9.37\times 10^{-3}$ $4.08\times 10^{-1}$ 1 $3.88\times 10^{-1}$ $8.75\times 10^{-1}$ $2.55\times 10^{-1}$ $5.54\times 10^{-2}$ $9.96\times 10^{-1}$ $6.51\times 10^{-1}$
2 $3.33\times 10^{-4}$ $3.10\times 10^{-1}$ 1 $3.99\times 10^{-1}$ $8.73\times 10^{-1}$ $3.15\times 10^{-1}$ $5.29\times 10^{-2}$ $9.96\times 10^{-1}$ $6.76\times 10^{-1}$
3 $4.65\times 10^{-3}$ $4.10\times 10^{-1}$ 1 $3.60\times 10^{-1}$ $8.64\times 10^{-1}$ $2.95\times 10^{-1}$ $6.67\times 10^{-2}$ $9.96\times 10^{-1}$ $6.61\times 10^{-1}$
4 $1.19\times 10^{-4}$ $3.93\times 10^{-1}$ 1 $4.02\times 10^{-1}$ $8.77\times 10^{-1}$ $3.15\times 10^{-1}$ $5.54\times 10^{-2}$ $9.96\times 10^{-1}$ $6.78\times 10^{-1}$
5 $1.53\times 10^{-4}$ $3.46\times 10^{-1}$ 1 $4.30\times 10^{-1}$ $8.82\times 10^{-1}$ $3.35\times 10^{-1}$ $5.29\times 10^{-2}$ $9.96\times 10^{-1}$ $6.45\times 10^{-1}$
$f_{24}$ 1 $2.46\times 10^{-1}$ $3.36\times 10^{-1}$ 1 $5.15\times 10^{-1}$ $8.52\times 10^{-1}$ $7.36\times 10^{-1}$ $2.78\times 10^{-2}$ $1.36\times 10^{0}$ $5.77\times 10^{-1}$
2 $2.91\times 10^{-1}$ $4.08\times 10^{-1}$ 2 $5.44\times 10^{-1}$ $8.59\times 10^{-1}$ $8.16\times 10^{-1}$ $2.01\times 10^{-2}$ $1.44\times 10^{0}$ $5.76\times 10^{-1}$
3 $2.51\times 10^{-1}$ $4.72\times 10^{-1}$ 2 $5.37\times 10^{-1}$ $8.60\times 10^{-1}$ $7.96\times 10^{-1}$ $1.92\times 10^{-2}$ $1.42\times 10^{0}$ $5.72\times 10^{-1}$
4 $2.29\times 10^{-1}$ $3.40\times 10^{-1}$ 1 $5.25\times 10^{-1}$ $8.48\times 10^{-1}$ $7.16\times 10^{-1}$ $3.04\times 10^{-2}$ $1.32\times 10^{0}$ $5.92\times 10^{-1}$
5 $2.54\times 10^{-1}$ $4.29\times 10^{-1}$ 2 $5.22\times 10^{-1}$ $8.46\times 10^{-1}$ $7.76\times 10^{-1}$ $2.20\times 10^{-2}$ $1.36\times 10^{0}$ $5.66\times 10^{-1}$
3 Explainablity of PSO
In this section, a novel explainability frame work is proposed for PSO. Mathematically, the following two equations govern the PSO framework:
$v_{i}(t+1)=wv_{i}(t)+c_{1}r_{1}(p_{best,i}-x_{i}(t))+c_{2}r_{2}(g_{best}-x_{i}(t)),$ (1)
$x_{i}(t+1)=x_{i}(t)+v_{i}(t+1).$ (2)
The movement dynamics of the PSO algorithm is governed by these two fundamental equations (1 and 2). Eq. (1) determines velocity, while Eq. (2) controls position of the particles (candidate solutions). In these formulations, $v_{i}(t)$ represents the velocity vector and $x_{i}(t)$ denotes the position vector of the $i^{th}$ particle in iteration $t$.
Apart from these parameter values, variations in communication topologies significantly influence interactions and information sharing among the particles, The neighbourhood topology in PSO is generally constructed with bidirectional connections, meaning particles maintain mutual influence if particle $j$ belongs to the neighbourhood particle $i$, then $i$ is reciprocally included in $j$’s neighbourhood. Within this structure, particles exchange information with their connected neighbours and adjust their trajectories based on the best personal solution $(pbest)$ discovered within their local network. While the algorithm theoretically permits numerous topological configurations, research and practice have predominantly focused on several established patterns.
3.1 Population Topologies
In this subsection, we are presenting some important communication topologies in PSO, The Star topology establishes a centralized communication framework where every particle in the swarm maintains a direct connection to the single best solution found by any member of the population. This architecture creates a fully-connected information network where all particles simultaneously receive updates from the global best position [37, 38]. The topology’s design ensures immediate propagation of the best-known solution throughout the entire swarm, creating strong directional guidance for all particles’ movement vectors. The structural configuration inherently prioritizes the dissemination of elite solutions while maintaining uniform influence across all swarm members. This complete connectivity pattern represents the most information-rich network configuration in swarm intelligence systems. The Ring topology organizes particles in a circular network where each individual maintains connections only to its immediate $k$ nearest neighbors. This decentralized architecture supports two operational modes: a static configuration with fixed neighborhood relationships, or a dynamic variant where connections may adapt during the search process. Particle interactions are governed by Minkowski $p$-norm distance metrics, with $p=1$ implementing Manhattan distance ($L_{1}$) and $p=2$ yielding Euclidean distance ($L_{2}$) calculations between swarm members. The constrained connectivity pattern creates localized information channels that naturally limit solution homogenization across the population [29, 47]. The Von Neumann topology organizes particles in a multidimensional grid configuration, where connectivity follows Delannoy number relationships. These combinatorial patterns determine each particle’s neighborhood size as a function of both the specified interaction range ($r$) and the problem’s dimensionality. The resulting lattice structure creates regular, spatially-distributed connections between particles, establishing a systematic information flow pattern across the search space. This geometric arrangement provides distinct advantages in maintaining organized particle interactions while preserving structured exploration capabilities [51, 34].
3.2 Explainer Framework for PSO
In this section, we propose an explainer framework for PSO built upon the IOHxplainer [50] environment. IOHxplainer provides a robust foundation for capturing, visualising, and quantifying algorithm dynamics by linking performance trajectories with exploratory landscape analysis (ELA) features and other descriptive measures. By extending this framework to PSO, we aim to uncover interpretable insights into swarm behaviour, parameter sensitivities, and landscape-driven performance variations. Such a framework will not only enhance transparency in understanding PSO but will also support principled algorithm selection, parameter tuning, and the design of adaptive strategies for complex optimization tasks.
3.2.1 IOHxplainer
IOHxplainer is a benchmarking platform designed to assess the performance of iterative optimization heuristics (IOHs), such as evolutionary algorithms and swarm-based algorithms [50]. It leverages explainable AI (XAI) to automatically generate insightful visualizations and statistical analyses, streamlining the evaluation of millions of algorithm configurations from large-scale empirical studies. This method helps researchers better understand how different algorithm parts work together and affect performance in various problem areas. The system collects performance data, including anytime performance measures, by testing many different algorithm setups on a range of benchmark functions. To use IOHxplainer, users start by defining the configuration space for the algorithm they want to study. This means setting value limits for continuous, integer, and categorical parameters and laying out any conditional relationships between them. The platform supports both thorough grid searches and random sampling, making it easier to explore and analyze potentially large configuration spaces. Although IOHxplainer is especially advantageous for examining modular algorithms with numerous configuration possibilities, it is also suitable for studying conventional algorithms along with their hyperparameter spaces. The configuration space is defined in a flexible manner, similar to approaches like SMAC  [15, 28], ensuring adaptability to different optimization scenarios.
Figure 1: A Framework for Automated and Explainable Analysis of PSO Configurations.
In this research, we utilized the IOHxplainer framework, originally designed and tested for evolutionary algorithms such as modCMA [26] and modDE [49]. We have re-invented this framework to analyze and elucidate the behavior of PSO algorithm, specifically to analyse the dynamics of PSO integrated with various communication topologies. A detailed representation of this framework is depicted in Figure 1.
The integration of PSO into the IOHxplainer framework followed a structured, three-stage experimental pipeline. First, the core PSO parameters were defined and fixed to establish a standardized configuration space, as detailed in Table 7. Second, this base algorithm was extended with three distinct communication topologies namely, Star, Ring, and Von Neumann, and each variant was rigorously evaluated across the diverse set of BBOB benchmark functions, recording performance based on function evaluations. Finally, the experimental data generation phase, labeled as AOCC performance data, involved the meticulous execution of the configured PSO algorithms across all BBOB functions to produce rich, time-series datasets. This process required running numerous independent repetitions for each algorithm-topology-function combination to ensure statistical robustness, all while strictly adhering to a fixed budget of function evaluations. Crucially, during every run, a detailed log was automatically generated, capturing the entire optimization trajectory by recording the best-so-far fitness value at a high frequency, often after every evaluation, thus creating a complete record of performance convergence over time. This resulted in a comprehensive “entire run data” output, a standardized collection of files containing the full history of each experiment, which was then perfectly structured for subsequent ingestion and analysis by the IOHprofiler tools. Algorithm 1 gives the detailed algorithm framework of integration of IOHxplainer with PSO.
Algorithm 1 Integrated PSO–IOHxplainer Framework
1:Import: pyswarms, ConfigurationSpace, Star, Ring, VonNeumann, explainer, seaborn
2:Define configuration space: confSpace: $\{c_{1},c_{2},w,n\_{\text{particles}},k,p,r\}$ with discrete parameter values
3:Initialize topology: Topology()
4:Define function run_pso(func, config, budget, dim):
5:   Extract hyperparameters from config
6:   Initialize PSO via GeneralOptimizerPSO
7:   return optimizer.optimize(func, iters=budget)
8:Initialize: explainer(run_pso, confSpace, algname="PSO")
9:Set experimental parameters: dimension,   func,   iids=[1,2,3,4,5],   reps=3
10:Configure sampling method: "grid", budget, seed=1
11:Execute: explainer.run(parallel=False, checkpoint="data.csv")
12:Save results: explainer.save_results("results.pkl")
13:Generate statistics: df = explainer.performance_stats()
14:Explain results: explainer.explain(partial_dependence=True, best_config=True)
These enhancements aim to clarify the effects of topology configurations on the algorithm’s performance by employing a suite of XAI techniques. This enables the extraction of insightful visualizations and statistical data from detailed empirical analyses which facilitates the evaluation of potentially innumerable PSO algorithmic frameworks. It also allows for the definition and consideration of hyper-parameter dependencies, providing a nuanced approach to the configuration of the PSO algorithm and its considered topologies. For each PSO configuration, we establish a specific evaluation budget and dimensionality to test against the BBOB benchmark suite [19], which features 24 diverse noiseless functions as part of the COCO framework [17].
Upon completing the experimental runs, which are executed in parallel, the framework captures detailed performance metrics based on pre-set budget, pre-set target, or flexible-time performance measures. The ‘explainer’ component of the framework then applies advanced XAI methodologies, such as the TreeSHAP technique [32], to determine the SHAP values for components of the PSO algorithm and associated topology hyperparameters  [33]. These values indicate the marginal impact of each element on the performance metrics across different runs. Although typically approximated due to computational constraints, the SHAP values are meticulously aggregated and visualized for each function within the benchmark suite. This visualization, often presented as a swarm plot, provides PSO designers and practitioners with a clear view of how specific hyperparameters influence performance under varied conditions and across different topologies. This data, once gathered, is processed either in combination with previous datasets or as a stand-alone analysis to extract SHAP values. The details of experiments and results are compiled in subsequent sections and related resources are provided at GitHub link: https://github.com/GitNitin02/ioh_pso.
4 Experimental Setup and Results
This section presents the experimental framework for evaluating PSO and its topological variants, with particular focus on explainability through our proposed analytical framework (Subsection 3.2.1). Our evaluation utilizes the comprehensive BBOB benchmark suite, which includes 24 distinct functions categorized by their mathematical properties and optimization challenges. The benchmark set comprises five classes: separable functions ($f_{1}$-$f_{5}$); functions with low-moderate conditioning ($f_{6}$ - $f_{9}$); highly conditioned, unimodal functions ($f_{10}$ - $f_{14}$); structured multimodal functions ($f_{15}$ - $f_{19}$); and weakly structured multimodal functions ($f_{20}$ - $f_{24}$). Within these groups, we further distinguish between unimodal ($f_{1}$, $f_{2}$, $f_{5}$ - $f_{14}$), multimodal ($f_{3}$, $f_{4}$, $f_{15}$, $f_{16}$, $f_{20}$, $f_{23}$, $f_{24}$), and highly multimodal ($f_{17}$ - $f_{19}$, $f_{21}$, $f_{22}$) functions. This carefully curated selection enables us to thoroughly assess PSO’s performance across different problem types, from simple convex landscapes to complex, rugged optimization surfaces. The combination of these diverse benchmark functions with our explainability framework provides both quantitative performance metrics and qualitative insights into how different PSO configurations behave under various optimization challenges. Through this experimental design, we can systematically evaluate the strengths and limitations of each topological variant across the entire spectrum of optimization difficulty.
Our experimental evaluation examines all twenty four BBOB functions across two problem dimensions (2d and 5d) for each PSO configuration. The computational budget is set at 100 and 500 iterations, with three distinct swarm topologies Star, Ring, and Von Neumann. The baseline PSO parameters for each topology in both dimensional spaces are detailed in Table 7. Our methodology for ensuring statistical reliability involves conducting five independent runs across the first five instances of each benchmark function. Algorithm performance is quantified using the normalized Area Over the Convergence Curve (AOCC) (Eq. (3)), which provides a comprehensive measure of optimization efficiency by integrating solution quality across the entire search process [50]. This metric enables direct comparison of different configurations’ effectiveness throughout the optimization trajectory.
Table 7: PSO module and their configurable hyperparameter for each topologies for d=2 & 5.
Hyperparameter Shorthand Domain
Cognitive coefficient $c_{1}$ {0.3, 0.5, 0.7, 0.9}
Social coefficient $c_{2}$ {0.2, 0.4, 0.6, 0.7}
Inertia weight $w$ {0.9, 0.5, 0.7}
Number of particles $n$ {50, 100, 150}
Nearest $k$ neighbors $k$ {1, 2, 3}
Minkowski $p$-norm $p$ {1, 2}
Delannoy numbers $r$ {1, 2}
$AOCC(\bar{y})=\frac{1}{B}\sum_{i=1}^{B}\left[1-{(\min(\max((y_{i}),lb),ub)-lb)}/{(ub-lb)}\right]$ (3)
AOCC metric is calculated using the series of best-found function values ($\bar{y}$) within a given evaluation budget ($B$), where the function value range is bounded between $lb$ (lower bound) and $ub$ (upper bound). This performance measure builds upon the empirical cumulative distribution function (ECDF), [31] which characterizes the probability distribution of optimization results. Specifically, AOCC represents the integrated area beneath the ECDF curve across an infinite set of target values within the specified bounds.
Following standard BBOB benchmarking practices, all function values were logarithmically transformed prior to the AOCC calculation, with values cut to the [5,5][-5,5] range for $2d$ and $5d$ problems. The experimental settings consisted of two computing platforms: $2d$ evaluations were performed on an Intel i7 system with 32GB RAM and 8 processing cores, while 5d experiments utilized a more powerful Intel Xeon W-1270P workstation with 96GB RAM and 8 cores to accommodate the increased computational demands of higher-dimensional optimization.
5 Hyperparameter Impact Analysis
We examined $1,728$ PSO configurations across twenty four BBOB functions, focusing on the impact of different communication topologies and hyperparameters on performance. Using SHAP-based plots for 2d & 5d (Table 7), we analyse the contribution parameters such as n_particles, $c_{1}$, $c_{2}$, $p$, and $k$. The SHAP values indicate how each parameter affects the AOCC, with yellow dots representing positive contributions and violet dots indicating negative ones. There is a clear distinction between the two colours across all topologies and parameters, such as n_particles, $c_{1}$, $c_{2}$, $p$, $k$, etc. and others. Specifically, yellow dots represent positive SHAP values, while violet dots consistently indicate negative SHAP values.
5.1 Experimental findings and Discussion for d=2
The distinct impact of topology on functions $f_{1}$, $f_{3}$, and $f_{17}$ is illustrated in Figures 2-4, for each Star, Ring, and Von Neumann architectures. When we observe that the impact of hyperparameter function $f_{1}$ (uni-model) changes depend on all the three considered topology in our study i.e., Star, Ring, or Von Neumann. In the Star setup, the social acceleration coefficient $c_{2}$ and the inertia weight $w$ play a big role. When $c_{2}$ values are higher, we see better performance because of a strong centralized effect. For the Ring topology, where information flows more gradually, the influences from $c_{1},c_{2}$, and $w$ are more modest, showing a more even balance without any one parameter taking over. In the Von Neumann arrangement, we can observe a wider variety of influences, especially from $c_{2}$ and $w$, which indicates that this grid-like structure helps with more adaptable convergence. Overall, the differences we see due to instance variance and randomness are quite small, meaning the effects of the hyper-parameters are consistent across various problem situations and are resilient against randomness in the PSO process. This points to a stable and trustworthy optimization method.
Function $f_{3}$ (multi-modal) exhibits strong and consistent performance across all three topologies, with minimal dependence on hyperparameter adjustments. In the Star topology, nearly all parameter contributions cluster near zero, reflecting stable performance regardless of configuration. The Ring topology follows the same pattern, displaying a flat distribution of contributions, further confirming its robustness to tuning. Similarly, the Von Neumann topology shows low variability, with no single parameter significantly influencing results. Additionally, both instance variance and stochastic variance remain consistently low in all topologies, demonstrating that $f_{3}$’s performance is highly reliable, unaffected by randomness or variations between problem instances. This makes it an excellent benchmark for PSO tuning in 2d optimization.
Function $f_{17}$ (highly multi-modal), in contrast, is highly sensitive to hyperparameter tuning across all topologies. In the Star topology, strong influences from $c_{1},c_{2}$, and $w$ indicate that convergence depends heavily on cognitive and social learning factors. The Ring topology shows slightly moderated but still notable contributions, particularly from $c_{2}$ and $w$, suggesting that information diffusion constraints partially but not entirely reduce sensitivity. The Von Neumann topology strikes a middle ground, with clear impacts from $c_{2},w$, and n_particles, emphasizing the role of swarm size and structure in balancing exploration. Despite this sensitivity, instance variance and stochastic variance remain low in all cases, meaning hyperparameter effects are consistent across different runs and problem instances. This ensures reliable and reproducible tuning outcomes, even for this more complex function.
Figure 2: The Impact of Hyperparameters on Performance for Benchmark Function in a 2d PSO with Star Topology
Figure 3: The Impact of Hyperparameters on Performance for Benchmark Function in a 2d PSO with Ring Topology
Figure 4: The Impact of Hyperparameters on Performance for Benchmark Function in a 2d PSO with Von Neumann Topology
5.2 Experimental findings and Discussion for d=5
Similarly, from Figures 5 -7 it can be observed that the functions $f_{10}$, $f_{3}$, and $f_{17}$ show significant variation in hyperparameter contributions across all three topologies, with $c_{2}$ emerging as the most influential. In the Star topology, contributions are relatively balanced but display greater variability in $c_{1},c_{2},$ and $k,$ highlighting heightened sensitivity to exploration and neighborhood size. The Ring topology, however, shifts this sensitivity more toward $c_{2}$, with $k$ playing a diminished role, suggesting a stronger reliance on local information. The Von Neumann topology also emphasizes $c_{2}$ and $r$ but with narrower distributions and reduced stochastic variance, indicating more consistent performance. While $c_{2}$ and $w$ remain key drivers across all topologies, the Von Neumann structure demonstrates the highest robustness, followed by Star and then Ring.
Similarly, for the functions $f_{3}$ and $f_{17}$, all topologies show clearer and more consistent parameter impacts. Star and Von Neumann have steady contributions from $w$, $c_{2}$ and $c_{1}$ enabling reliable convergence. Ring shows slightly more dispersion but still maintains stable influence, reflecting the easier landscape where local communication suffices.The n_particles parameter (number of particles in the swarm) demonstrates negative contributions (violet in SHAP plots) across all three topologies (Star, Ring, Von Neumann), indicating that increasing swarm size often degrades optimization performance. The repository and supplementary file now include the remaining graphs and plots for each function, providing a complete visualization of their behavior and outputs. The supplementary file is available at https://github.com/GitNitin02/ioh_pso/blob/main/Supplementary.pdf.
Figure 5: The Impact of Hyperparameters on Performance for Benchmark Function in a 5d PSO with Star Topology
Figure 6: The Impact of Hyperparameters on Performance for Benchmark Function in a 5d PSO with Ring Topology
Figure 7: The Impact of Hyperparameters on Performance for Benchmark Function in a 5d PSO with Von Neumann Topology
5.3 PSO Variant configuration selection
The “single-best mean” (sbm) and “single-best std” (sbs) are the statistical measures used to evaluate the consistency and performance of the global best solution ($g_{best}$) over multiple independent runs. The “sbm” represents the average of the best solutions found across all runs, indicating the algorithm’s overall effectiveness in locating near-optimal solutions. The “sbs” measures the variability in these solutions a low standard deviation suggests reliable convergence, while a high deviation indicates instability or sensitivity to initial conditions. Similarly, “avg-best mean” (abm) and “avg-best std” (abs) assess the collective behaviour of the swarm by analysing the average of particles’ personal best positions ($p_{best,i}$) over multiple runs. The abm reflects the typical performance level of the swarm, while the abs quantifies how consistently the swarm performs across different runs. A low abs implies stable exploration, whereas a high value suggests erratic convergence or poor swarm coordination. Together, these metrics help fine-tune PSO parameters, ensuring robust optimization by balancing exploration and exploitation.
[…]
Similarly, the results in Table 9 For unimodal functions, such as $f_{12}$, all topologies perform exceptionally well, achieving near-optimal results (e.g., means of $2.00\times 10^{-3}$ ) with negligible variability (standard deviations as low as $8.90\times 10^{-19}$), indicating robustness in simple landscapes. However, in multi-modal functions like $f_{4}$ and $f_{6}$, differences emerge. The Ring topology often excels in maintaining diversity, as seen in $f_{4}$ with a lower single-best standard deviation ($6.21\times 10^{-3}$) compared to the Star topology ($1.00\times 10^{-2}$). Meanwhile, the Von Neumann topology strikes a balance between exploration and exploitation, performing well in functions like $f_{6}$ with a high $R^{2}$ ratio ($9.70\times 10^{-1}$), significantly outperforming the Star topology ($8.84\times 10^{-3}$). For highly complex or deceptive functions, such as $f_{7}$ and $f_{14}$, the Von Neumann topology demonstrates greater resilience. For instance, in $f_{14}$, it achieves an overall mean of $1.81\times 10^{-2}$, while the Ring topology occasionally dominates in specific cases like $f_{7}$ with a $R^{2}$ ratio of $9.12\times 10^{-1}$. The Star topology, though fast in convergence, exhibits higher variability and inconsistency, as seen in $f_{7}$ with a single-best standard deviation of $2.75\times 10^{-2}$ and in $f_{9}$ with erratic performance despite a low overall mean ($1.00\times 10^{-2}$ ).
In summary, the choice of PSO topology should be guided by problem complexity. The Star topology is suitable for simple, unimodal problems due to its rapid convergence, while the Ring topology is preferable for rugged, multi-modal landscapes where diversity preservation is crucial. The Von Neumann topology, with its balanced approach, proves most effective for deceptive or highly complex functions. For real-world applications where problem characteristics are unknown a priori, adaptive or hybrid approaches may offer the best performance.
In the context of the PSO topology comparison, $R^{2}$ (R-squared) [36] could serve as a supplementary metric to evaluate how well each topology’s performance aligns with an ideal optimization trajectory or expected convergence behaviour. For example, if the Star topology’s fast convergence in unimodal functions consistently follows a predictable pattern, its $R^{2}$ value would be high, indicating that its performance is tightly correlated with the problem’s simplicity. Conversely, in multi-modal functions, a lower $R^{2}$ for the Star topology might reflect its erratic performance due to premature convergence, whereas the Ring or Von Neumann topologies with their better diversity maintenance could exhibit higher $R^{2}$ values, suggesting more stable alignment with the problem’s complexity.
5.3.1 Time Complexity Analysis
The impact of swarm topology on PSO’s computational efficiency becomes increasingly significant when examining different problem dimensions. For 2-dimensional problems, our experiments reveal clear differences in execution times: the Von Neumann topology demonstrates superior efficiency (6.56 hours), followed by Star (7.2 hours) and Ring (9.35 hours). This pattern persists and amplifies in 5-dimensional problems, with Von Neumann maintaining its advantage (56.87 hours), while Star and Ring require 61.17 hours and 75.52 hours, respectively. The consistent performance gap across dimensions suggests that Von Neumann’s structured yet decentralized communication pattern scales more effectively, likely due to its balanced information flow that avoids bottlenecks. In contrast, the Ring topology’s sequential neighbor-based updates appear to compound its time complexity as dimensionality increases, resulting in significantly longer runtimes. The Star topology occupies an intermediate position, with its global communication structure offering better scalability than Ring but still incurring higher costs than Von Neumann. These findings underscore how topology selection becomes increasingly critical for computational efficiency in higher-dimensional optimization problems, where the differences in time complexity grow substantially. The results provide practical guidance for algorithm design, particularly in scenarios where both solution quality and computational resources must be carefully balanced.
Table 8: Performance of Star, Ring and Von Neumann Topologies over all BBOB functions on d=2.
Function Topologies sbm sbs abm abs all mean all std R2 train
$f_{1}$ Star $2.50\times 10^{-1}$ $5.40\times 10^{-2}$ $2.23\times 10^{-1}$ $2.62\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.34\times 10^{-1}}$ $4.18\times 10^{-2}$ $9.76\times 10^{-1}$
Ring $2.50\times 10^{-1}$ $5.40\times 10^{-2}$ $2.23\times 10^{-1}$ $2.62\times 10^{-2}$ $2.32\times 10^{-1}$ $4.32\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.83\times 10^{-1}}$
Von Neumann $2.50\times 10^{-1}$ $5.40\times 10^{-2}$ $2.23\times 10^{-1}$ $2.62\times 10^{-2}$ $2.33\times 10^{-1}$ $4.25\times 10^{-2}$ $9.80\times 10^{-1}$
$f_{2}$ Star $2.35\times 10^{-2}$ $2.32\times 10^{-2}$ $1.70\times 10^{-2}$ $1.89\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.86\times 10^{-2}}$ $2.23\times 10^{-2}$ $9.86\times 10^{-1}$
Ring $2.39\times 10^{-2}$ $2.67\times 10^{-2}$ $1.70\times 10^{-2}$ $1.89\times 10^{-2}$ $1.84\times 10^{-2}$ $2.23\times 10^{-2}$ $9.87\times 10^{-1}$
Von Neumann $2.39\times 10^{-2}$ $2.67\times 10^{-2}$ $1.70\times 10^{-2}$ $1.89\times 10^{-2}$ $1.85\times 10^{-2}$ $2.22\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.88\times 10^{-1}}$
$f_{3}$ Star $1.01\times 10^{-1}$ $2.82\times 10^{-2}$ $1.01\times 10^{-1}$ $2.82\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.32\times 10^{-2}}$ $2.09\times 10^{-2}$ $9.66\times 10^{-1}$
Ring $1.01\times 10^{-1}$ $2.82\times 10^{-2}$ $1.01\times 10^{-1}$ $2.82\times 10^{-2}$ $9.27\times 10^{-2}$ $2.12\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.77\times 10^{-1}}$
Von Neumann $1.01\times 10^{-1}$ $2.82\times 10^{-2}$ $1.01\times 10^{-1}$ $2.82\times 10^{-2}$ $9.30\times 10^{-2}$ $2.11\times 10^{-2}$ $9.69\times 10^{-1}$
$f_{4}$ Star $9.62\times 10^{-2}$ $2.99\times 10^{-2}$ $7.81\times 10^{-2}$ $1.45\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}8.68\times 10^{-2}}$ $2.23\times 10^{-2}$ $9.87\times 10^{-1}$
Ring $9.51\times 10^{-2}$ $3.13\times 10^{-2}$ $7.81\times 10^{-2}$ $1.45\times 10^{-2}$ $8.62\times 10^{-2}$ $2.26\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.93\times 10^{-1}}$
Von Neumann $9.87\times 10^{-2}$ $3.06\times 10^{-2}$ $7.81\times 10^{-2}$ $1.45\times 10^{-2}$ $8.66\times 10^{-2}$ $2.24\times 10^{-2}$ $9.83\times 10^{-1}$
$f_{5}$ Star $1.55\times 10^{-1}$ $2.72\times 10^{-2}$ $1.55\times 10^{-1}$ $2.72\times 10^{-2}$ $1.37\times 10^{-1}$ $2.33\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.96\times 10^{-1}}$
Ring $1.55\times 10^{-1}$ $2.72\times 10^{-2}$ $1.55\times 10^{-1}$ $2.72\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.38\times 10^{-1}}$ $2.26\times 10^{-2}$ $9.94\times 10^{-1}$
Von Neumann $1.55\times 10^{-1}$ $2.72\times 10^{-2}$ $1.55\times 10^{-1}$ $2.72\times 10^{-2}$ $1.37\times 10^{-1}$ $2.32\times 10^{-2}$ $9.90\times 10^{-1}$
$f_{6}$ Star $1.41\times 10^{-1}$ $3.49\times 10^{-2}$ $1.38\times 10^{-1}$ $2.33\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.33\times 10^{-1}}$ $2.85\times 10^{-2}$ $9.79\times 10^{-1}$
Ring $1.38\times 10^{-1}$ $2.33\times 10^{-2}$ $1.38\times 10^{-1}$ $2.33\times 10^{-2}$ $1.31\times 10^{-1}$ $2.92\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.94\times 10^{-1}}$
Von Neumann $1.44\times 10^{-1}$ $3.46\times 10^{-2}$ $1.38\times 10^{-1}$ $2.33\times 10^{-2}$ $1.32\times 10^{-1}$ $2.89\times 10^{-2}$ $9.78\times 10^{-1}$
$f_{7}$ Star $1.92\times 10^{-1}$ $3.16\times 10^{-2}$ $1.78\times 10^{-1}$ $4.26\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.84\times 10^{-1}}$ $3.81\times 10^{-2}$ $9.64\times 10^{-1}$
Ring $1.91\times 10^{-1}$ $4.60\times 10^{-2}$ $1.78\times 10^{-1}$ $4.26\times 10^{-2}$ $1.83\times 10^{-1}$ $3.81\times 10^{-2}$ $9.80\times 10^{-1}$
Von Neumann $1.91\times 10^{-1}$ $4.60\times 10^{-2}$ $1.78\times 10^{-1}$ $4.26\times 10^{-2}$ $1.83\times 10^{-1}$ $3.80\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.82\times 10^{-1}}$
$f_{8}$ Star $1.46\times 10^{-1}$ $4.30\times 10^{-2}$ $1.33\times 10^{-1}$ $5.85\times 10^{-2}$ $1.34\times 10^{-1}$ $4.82\times 10^{-2}$ $9.69\times 10^{-1}$
Ring $1.42\times 10^{-1}$ $4.32\times 10^{-2}$ $1.33\times 10^{-1}$ $5.85\times 10^{-2}$ $1.34\times 10^{-1}$ $4.80\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.84\times 10^{-1}}$
Von Neumann $1.44\times 10^{-1}$ $3.82\times 10^{-2}$ $1.33\times 10^{-1}$ $5.85\times 10^{-2}$ $1.34\times 10^{-1}$ $4.82\times 10^{-2}$ $9.77\times 10^{-1}$
$f_{9}$ Star $1.63\times 10^{-1}$ $3.81\times 10^{-2}$ $1.63\times 10^{-1}$ $3.81\times 10^{-2}$ $1.43\times 10^{-1}$ $4.74\times 10^{-2}$ $9.54\times 10^{-1}$
Ring $1.63\times 10^{-1}$ $3.81\times 10^{-2}$ $1.63\times 10^{-1}$ $3.81\times 10^{-2}$ $1.42\times 10^{-1}$ $4.74\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.88\times 10^{-1}}$
Von Neumann $1.63\times 10^{-1}$ $3.81\times 10^{-2}$ $1.63\times 10^{-1}$ $3.81\times 10^{-2}$ $1.43\times 10^{-1}$ $4.88\times 10^{-2}$ $9.63\times 10^{-1}$
$f_{10}$ Star $2.36\times 10^{-2}$ $3.04\times 10^{-2}$ $2.36\times 10^{-2}$ $3.04\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.60\times 10^{-2}}$ $1.85\times 10^{-2}$ $9.61\times 10^{-1}$
Ring $2.36\times 10^{-2}$ $3.04\times 10^{-2}$ $2.36\times 10^{-2}$ $3.04\times 10^{-2}$ $1.54\times 10^{-2}$ $1.83\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.82\times 10^{-1}}$
Von Neumann $2.36\times 10^{-2}$ $3.04\times 10^{-2}$ $2.36\times 10^{-2}$ $3.04\times 10^{-2}$ $1.57\times 10^{-2}$ $1.85\times 10^{-2}$ $9.54\times 10^{-1}$
$f_{11}$ Star $6.00\times 10^{-2}$ $4.22\times 10^{-2}$ $6.00\times 10^{-2}$ $4.22\times 10^{-2}$ $6.00\times 10^{-2}$ $4.01\times 10^{-2}$ $9.80\times 10^{-1}$
Ring $6.00\times 10^{-2}$ $4.22\times 10^{-2}$ $6.00\times 10^{-2}$ $4.22\times 10^{-2}$ $6.00\times 10^{-2}$ $4.01\times 10^{-2}$ $9.80\times 10^{-1}$
Von Neumann $5.90\times 10^{-2}$ $4.22\times 10^{-2}$ $6.00\times 10^{-2}$ $4.22\times 10^{-2}$ $6.00\times 10^{-2}$ $4.01\times 10^{-2}$ $9.77\times 10^{-1}$
$f_{12}$ Star $1.73\times 10^{-2}$ $1.78\times 10^{-2}$ $1.53\times 10^{-2}$ $1.28\times 10^{-2}$ $1.35\times 10^{-2}$ $1.31\times 10^{-2}$ $9.43\times 10^{-1}$
Ring $1.62\times 10^{-2}$ $1.30\times 10^{-2}$ $1.44\times 10^{-2}$ $1.25\times 10^{-2}$ $1.35\times 10^{-2}$ $1.26\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.91\times 10^{-1}}$
Von Neumann $1.84\times 10^{-2}$ $1.53\times 10^{-2}$ $1.84\times 10^{-2}$ $1.53\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.36\times 10^{-1}}$ $1.28\times 10^{-2}$ $9.58\times 10^{-1}$
$f_{13}$ Star $7.27\times 10^{-2}$ $2.37\times 10^{-2}$ $6.76\times 10^{-2}$ $2.33\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}6.27\times 10^{-2}}$ $2.72\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.72\times 10^{-1}}$
Ring $7.22\times 10^{-2}$ $2.47\times 10^{-2}$ $6.94\times 10^{-2}$ $2.60\times 10^{-2}$ $6.22\times 10^{-2}$ $2.74\times 10^{-2}$ $9.58\times 10^{-1}$
Von Neumann $7.19\times 10^{-2}$ $2.54\times 10^{-2}$ $6.62\times 10^{-2}$ $2.39\times 10^{-2}$ $6.19\times 10^{-2}$ $2.76\times 10^{-2}$ $9.70\times 10^{-1}$
$f_{14}$ Star $2.51\times 10^{-1}$ $3.35\times 10^{-2}$ $2.47\times 10^{-1}$ $3.40\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.40\times 10^{-1}}$ $3.39\times 10^{-2}$ $9.75\times 10^{-1}$
Ring $2.43\times 10^{-1}$ $3.72\times 10^{-2}$ $2.38\times 10^{-1}$ $3.67\times 10^{-2}$ $2.38\times 10^{-1}$ $3.45\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.85\times 10^{-1}}$
Von Neumann $2.49\times 10^{-1}$ $3.32\times 10^{-2}$ $2.46\times 10^{-1}$ $3.69\times 10^{-2}$ $2.39\times 10^{-1}$ $3.43\times 10^{-2}$ $9.66\times 10^{-1}$
$f_{15}$ Star $1.09\times 10^{-1}$ $3.02\times 10^{-2}$ $1.05\times 10^{-1}$ $3.30\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.01\times 10^{-1}}$ $2.40\times 10^{-2}$ $9.82\times 10^{-1}$
Ring $1.07\times 10^{-1}$ $3.14\times 10^{-2}$ $1.04\times 10^{-1}$ $3.35\times 10^{-2}$ $1.00\times 10^{-1}$ $2.41\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.85\times 10^{-1}}$
Von Neumann $1.08\times 10^{-1}$ $3.13\times 10^{-2}$ $1.06\times 10^{-1}$ $3.19\times 10^{-2}$ $1.00\times 10^{-1}$ $2.41\times 10^{-2}$ $9.84\times 10^{-1}$
$f_{16}$ Star $1.68\times 10^{-1}$ $5.52\times 10^{-2}$ $1.48\times 10^{-1}$ $3.17\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.58\times 10^{-1}}$ $3.86\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.81\times 10^{-1}}$
Ring $1.68\times 10^{-1}$ $5.52\times 10^{-2}$ $1.47\times 10^{-1}$ $3.16\times 10^{-2}$ $1.57\times 10^{-1}$ $3.83\times 10^{-2}$ $9.80\times 10^{-1}$
Von Neumann $1.68\times 10^{-1}$ $5.52\times 10^{-2}$ $1.54\times 10^{-1}$ $3.65\times 10^{-2}$ $1.57\times 10^{-1}$ $3.87\times 10^{-2}$ $9.74\times 10^{-1}$
$f_{17}$ Star $1.74\times 10^{-1}$ $1.32\times 10^{-2}$ $1.74\times 10^{-1}$ $1.32\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.68\times 10^{-1}}$ $1.74\times 10^{-2}$ $9.62\times 10^{-1}$
Ring $1.73\times 10^{-1}$ $1.59\times 10^{-2}$ $1.71\times 10^{-2}$ $1.50\times 10^{-2}$ $1.67\times 10^{-1}$ $1.72\times 10^{-1}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.82\times 10^{-1}}$
Von Neumann $1.75\times 10^{-1}$ $1.91\times 10^{-2}$ $1.74\times 10^{-1}$ $1.96\times 10^{-2}$ $1.67\times 10^{-1}$ $1.76\times 10^{-2}$ $9.53\times 10^{-1}$
$f_{18}$ Star $1.25\times 10^{-1}$ $2.17\times 10^{-2}$ $1.25\times 10^{-1}$ $55\times 10^{-2} 1.1$$11\times 10^{-1}$ $2.27\times 10^{-2}$ $9.68\times 10^{-1}$
Ring $1.23\times 10^{-1}$ $1.89\times 10^{-2}$ $1.21\times 10^{-1}$ $2.14\times 10^{-2}$ $1.11\times 10^{-1}$ $2.29\times 10^{-2}$ $9.68\times 10^{-1}$
Von Neumann $1.27\times 10^{-1}$ $1.78\times 10^{-2}$ $1.21\times 10^{-1}$ $87\times 10^{-2} 1.1$$11\times 10^{-1}$ $2.27\times 10^{-2}$ $9.69\times 10^{-1}$
$f_{19}$ Star $2.45\times 10^{-1}$ $5.90\times 10^{-2}$ $2.34\times 10^{-1}$ $5.59\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.21\times 10^{-1}}$ $4.71\times 10^{-2}$ $9.56\times 10^{-1}$
Ring $2.37\times 10^{-1}$ $5.41\times 10^{-2}$ $2.31\times 10^{-1}$ $5.96\times 10^{-2}$ $2.19\times 10^{-1}$ $4.75\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.82\times 10^{-1}}$
Von Neumann $2.45\times 10^{-1}$ $5.71\times 10^{-2}$ $2.37\times 10^{-1}$ $5.32\times 10^{-2}$ $2.20\times 10^{-1}$ $4.72\times 10^{-2}$ $9.72\times 10^{-1}$
$f_{20}$ Star $1.48\times 10^{-1}$ $1.77\times 10^{-2}$ $1.41\times 10^{-1}$ $2.33\times 10^{-2}$ $1.44\times 10^{-1}$ $1.98\times 10^{-2}$ $9.95\times 10^{-1}$
Ring $1.48\times 10^{-1}$ $1.77\times 10^{-2}$ $1.41\times 10^{-1}$ $2.49\times 10^{-2}$ $1.44\times 10^{-1}$ $2.01\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.98\times 10^{-1}}$
Von Neumann $1.51\times 10^{-1}$ $2.38\times 10^{-2}$ $1.50\times 10^{-1}$ $2.94\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1.45\times 10^{-1}}$ $2.04\times 10^{-2}$ $9.92\times 10^{-1}$
$f_{21}$ Star $2.37\times 10^{-1}$ $6.59\times 10^{-2}$ $2.22\times 10^{-1}$ $4.00\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.21\times 10^{-1}}$ $4.76\times 10^{-2}$ $9.62\times 10^{-1}$
Ring $2.31\times 10^{-1}$ $5.80\times 10^{-2}$ $2.21\times 10^{-1}$ $3.89\times 10^{-2}$ $2.19\times 10^{-1}$ $4.51\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.89\times 10^{-1}}$
Von Neumann $2.37\times 10^{-1}$ $6.57\times 10^{-2}$ $2.21\times 10^{-1}$ $3.48\times 10^{-2}$ $2.19\times 10^{-1}$ $4.58\times 10^{-2}$ $9.57\times 10^{-1}$
$f_{22}$ Star $2.37\times 10^{-1}$ $7.91\times 10^{-2}$ $2.37\times 10^{-1}$ $7.91\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2.20\times 10^{-1}}$ $5.78\times 10^{-2}$ $9.65\times 10^{-1}$
Ring $2.26\times 10^{-1}$ $6.61\times 10^{-2}$ $2.19\times 10^{-1}$ $6.01\times 10^{-2}$ $2.19\times 10^{-1}$ $6.03\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.72\times 10^{-1}}$
Von Neumann $2.27\times 10^{-1}$ $8.41\times 10^{-2}$ $2.22\times 10^{-1}$ $4.87\times 10^{-2}$ $2.19\times 10^{-1}$ $5.76\times 10^{-2}$ $9.64\times 10^{-1}$
$f_{23}$ Star $1.49\times 10^{-1}$ $2.22\times 10^{-2}$ $1.39\times 10^{-1}$ $2.57\times 10^{-2}$ $1.41\times 10^{-1}$ $2.13\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.78\times 10^{-1}}$
Ring $1.49\times 10^{-1}$ $2.22\times 10^{-2}$ $1.40\times 10^{-1}$ $2.49\times 10^{-2}$ $1.42\times 10^{-1}$ $2.11\times 10^{-2}$ $9.76\times 10^{-1}$
Von Neumann $1.49\times 10^{-1}$ $2.22\times 10^{-2}$ $1.44\times 10^{-1}$ $2.22\times 10^{-2}$ $1.42\times 10^{-1}$ $2.12\times 10^{-2}$ $9.66\times 10^{-1}$
$f_{24}$ Star $1.23\times 10^{-1}$ $60\times 10^{-2} 1.1$$14\times 10^{-1}$ $26\times 10^{-2} 1.1$$8\times 10^{-1} 1.1$$18\times 10^{-1}$ $9.52\times 10^{-1}$
Ring $1.23\times 10^{-1}$ $1.60\times 10^{-2}$ $1.33\times 10^{-2}$ $26\times 10^{-2} 1.1$$17\times 10^{-1}$ $1.29\times 10^{-2}$ $\mathbf{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9.66\times 10^{-1}}$
Von Neumann $1.23\times 10^{-1}$ $60\times 10^{-2} 1.1$$6\times 10^{-1} 1.1$$3\times 10^{-2} 1.1$$18\times 10^{-1}$ $1.26\times 10^{-2}$ $9.30\times 10^{-1}$
Table 9: Performance of Star, Ring and Von Neumann Topologies over all BBOB functions on $d=5$.
Function Topologies sbm sbs abm abs all mean all std R2 train
$f_{1}$ Star $2.20\times 10^{-1}$ $3.20\times 10^{-2}$ $2.20\times 10^{-1}$ $3.20\times 10^{-2}$ $1.54\times 10^{-1}$ $3.32\times 10^{-2}$ $9.32\times 10^{-1}$
Ring $1.37\times 10^{-1}$ $2.03\times 10^{-2}$ $1.37\times 10^{-1}$ $2.03\times 10^{-2}$ $1.18\times 10^{-1}$ $2.19\times 10^{-2}$ $8.95\times 10^{-1}$
Von Neumann $1.88\times 10^{-1}$ $2.88\times 10^{-2}$ $1.88\times 10^{-1}$ $2.88\times 10^{-2}$ $1.38\times 10^{-1}$ $2.79\times 10^{-2}$ $9.17\times 10^{-1}$
$f_{2}$ Star $2.91\times 10^{-3}$ $3.52\times 10^{-3}$ $2.91\times 10^{-3}$ $3.52\times 10^{-3}$ $2.02\times 10^{-3}$ $3.74\times 10^{-4}$ $2.83\times 10^{-1}$
Ring $2.00\times 10^{-3}$ $8.98\times 10^{-19}$ $2.00\times 10^{-3}$ $8.98\times 10^{-19}$ $2.00\times 10^{-3}$ $8.67\times 10^{-19}$ $3.27\times 10^{-1}$
Von Neumann $2.11\times 10^{-3}$ $4.29\times 10^{-4}$ $2.00\times 10^{-3}$ $8.98\times 10^{-19}$ $2.00\times 10^{-3}$ $4.28\times 10^{-5}$ $4.09\times 10^{-1}$
$f_{3}$ Star $5.00\times 10^{-2}$ $7.67\times 10^{-3}$ $5.00\times 10^{-2}$ $7.67\times 10^{-3}$ $3.57\times 10^{-2}$ $1.02\times 10^{-2}$ $7.93\times 10^{-1}$
Ring $3.28\times 10^{-2}$ $1.21\times 10^{-2}$ $3.28\times 10^{-2}$ $1.21\times 10^{-2}$ $2.38\times 10^{-2}$ $9.57\times 10^{-3}$ $8.00\times 10^{-1}$
Von Neumann $4.77\times 10^{-2}$ $9.53\times 10^{-3}$ $4.65\times 10^{-2}$ $8.14\times 10^{-3}$ $3.16\times 10^{-2}$ $9.71\times 10^{-3}$ $8.14\times 10^{-1}$
$f_{4}$ Star $4.00\times 10^{-2}$ $1.00\times 10^{-2}$ $4.00\times 10^{-2}$ $1.00\times 10^{-2}$ $2.64\times 10^{-2}$ $1.05\times 10^{-2}$ $8.00\times 10^{-1}$
Ring $2.14\times 10^{-2}$ $6.37\times 10^{-3}$ $2.05\times 10^{-2}$ $5.40\times 10^{-3}$ $1.18\times 10^{-2}$ $9.11\times 10^{-3}$ $8.40\times 10^{-1}$
Von Neumann $3.39\times 10^{-2}$ $1.05\times 10^{-2}$ $3.25\times 10^{-2}$ $9.29\times 10^{-3}$ $2.18\times 10^{-2}$ $1.00\times 10^{-2}$ $8.48\times 10^{-1}$
$f_{5}$ Star $7.15\times 10^{-2}$ $1.11\times 10^{-2}$ $7.15\times 10^{-2}$ $1.11\times 10^{-2}$ $6.22\times 10^{-2}$ $1.04\times 10^{-2}$ $8.85\times 10^{-1}$
Ring $6.48\times 10^{-2}$ $1.02\times 10^{-2}$ $6.24\times 10^{-2}$ $8.13\times 10^{-3}$ $5.84\times 10^{-2}$ $9.93\times 10^{-3}$ $9.27\times 10^{-1}$
Von Neumann $6.80\times 10^{-2}$ $7.90\times 10^{-3}$ $6.73\times 10^{-2}$ $7.79\times 10^{-3}$ $6.00\times 10^{-2}$ $9.58\times 10^{-3}$ $9.23\times 10^{-1}$
$f_{6}$ Star $7.40\times 10^{-2}$ $3.72\times 10^{-2}$ $7.40\times 10^{-2}$ $3.72\times 10^{-2}$ $4.06\times 10^{-2}$ $2.83\times 10^{-2}$ $8.84\times 10^{-1}$
Ring $4.30\times 10^{-2}$ $2.97\times 10^{-2}$ $4.27\times 10^{-2}$ $3.33\times 10^{-2}$ $2.16\times 10^{-2}$ $2.31\times 10^{-2}$ $8.81\times 10^{-1}$
Von Neumann $6.63\times 10^{-2}$ $2.94\times 10^{-2}$ $5.66\times 10^{-2}$ $2.82\times 10^{-2}$ $3.20\times 10^{-2}$ $2.59\times 10^{-2}$ $9.07\times 10^{-1}$
$f_{7}$ Star $1.30\times 10^{-1}$ $2.89\times 10^{-2}$ $1.30\times 10^{-1}$ $2.89\times 10^{-2}$ $9.84\times 10^{-2}$ $2.98\times 10^{-2}$ $8.81\times 10^{-1}$
Ring $9.81\times 10^{-2}$ $2.75\times 10^{-2}$ $9.24\times 10^{-2}$ $2.21\times 10^{-2}$ $7.08\times 10^{-2}$ $2.95\times 10^{-2}$ $9.12\times 10^{-1}$
Von Neumann $1.26\times 10^{-1}$ $1.80\times 10^{-2}$ $1.26\times 10^{-1}$ $1.80\times 10^{-2}$ $9.10\times 10^{-2}$ $2.91\times 10^{-2}$ $8.99\times 10^{-1}$
$f_{8}$ Star $5.20\times 10^{-2}$ $2.26\times 10^{-2}$ $5.20\times 10^{-2}$ $2.26\times 10^{-2}$ $1.30\times 10^{-2}$ $1.73\times 10^{-2}$ $8.17\times 10^{-1}$
Ring $1.09\times 10^{-2}$ $1.89\times 10^{-2}$ $1.09\times 10^{-2}$ $1.89\times 10^{-2}$ $3.41\times 10^{-3}$ $6.97\times 10^{-3}$ $8.63\times 10^{-1}$
Von Neumann $3.10\times 10^{-2}$ $1.36\times 10^{-2}$ $3.05\times 10^{-2}$ $1.40\times 10^{-2}$ $8.01\times 10^{-3}$ $1.16\times 10^{-2}$ $8.48\times 10^{-1}$
$f_{9}$ Star $4.78\times 10^{-2}$ $1.63\times 10^{-2}$ $4.78\times 10^{-2}$ $1.63\times 10^{-2}$ $1.49\times 10^{-2}$ $1.70\times 10^{-2}$ $8.41\times 10^{-1}$
Ring $6.53\times 10^{-3}$ $1.21\times 10^{-2}$ $6.53\times 10^{-3}$ $1.21\times 10^{-2}$ $2.98\times 10^{-3}$ $5.03\times 10^{-3}$ $9.19\times 10^{-1}$
Von Neumann $4.68\times 10^{-2}$ $1.12\times 10^{-2}$ $4.47\times 10^{-2}$ $1.34\times 10^{-2}$ $1.25\times 10^{-2}$ $1.54\times 10^{-2}$ $8.65\times 10^{-1}$
$f_{10}$ Star $2.12\times 10^{-3}$ $4.73\times 10^{-4}$ $2.12\times 10^{-3}$ $4.73\times 10^{-4}$ $2.00\times 10^{-3}$ $6.39\times 10^{-5}$ $3.49\times 10^{-1}$
Ring $2.30\times 10^{-3}$ $1.15\times 10^{-3}$ $2.00\times 10^{-3}$ $8.98\times 10^{-19}$ $2.00\times 10^{-3}$ $3.92\times 10^{-5}$ $4.17\times 10^{-1}$
Von Neumann $2.03\times 10^{-3}$ $1.33\times 10^{-4}$ $2.00\times 10^{-3}$ $4.41\times 10^{-19}$ $2.00\times 10^{-3}$ $1.38\times 10^{-5}$ $4.13\times 10^{-1}$
$f_{11}$ Star $3.87\times 10^{-2}$ $2.45\times 10^{-2}$ $3.87\times 10^{-2}$ $2.45\times 10^{-2}$ $1.79\times 10^{-2}$ $1.83\times 10^{-2}$ $6.58\times 10^{-1}$
Ring $3.49\times 10^{-2}$ $2.10\times 10^{-2}$ $3.49\times 10^{-2}$ $2.10\times 10^{-2}$ $1.47\times 10^{-2}$ $1.56\times 10^{-2}$ $6.54\times 10^{-1}$
Von Neumann $3.87\times 10^{-2}$ $2.40\times 10^{-2}$ $3.73\times 10^{-2}$ $2.01\times 10^{-2}$ $1.87\times 10^{-2}$ $1.76\times 10^{-2}$ $6.04\times 10^{-1}$
$f_{12}$ Star $2.00\times 10^{-3}$ $8.98\times 10^{-19}$ $2.00\times 10^{-3}$ $8.98\times 10^{-19}$ $2.00\times 10^{-3}$ $8.67\times 10^{-19}$ $7.34\times 10^{-1}$
Ring $2.00\times 10^{-3}$ $8.98\times 10^{-19}$ $2.00\times 10^{-3}$ $8.98\times 10^{-19}$ $2.00\times 10^{-3}$ $8.67\times 10^{-19}$ $7.34\times 10^{-1}$
Von Neumann $2.00\times 10^{-3}$ $4.41\times 10^{-19}$ $2.00\times 10^{-3}$ $4.41\times 10^{-19}$ $2.00\times 10^{-3}$ $8.67\times 10^{-19}$ $6.49\times 10^{-1}$
$f_{13}$ Star $1.25\times 10^{-2}$ $1.05\times 10^{-2}$ $1.25\times 10^{-2}$ $1.05\times 10^{-2}$ $3.63\times 10^{-3}$ $4.34\times 10^{-3}$ $6.81\times 10^{-1}$
Ring $3.94\times 10^{-3}$ $7.51\times 10^{-3}$ $2.00\times 10^{-3}$ $8.98\times 10^{-19}$ $2.06\times 10^{-3}$ $6.77\times 10^{-4}$ $7.11\times 10^{-1}$
Von Neumann $1.04\times 10^{-2}$ $7.41\times 10^{-3}$ $1.04\times 10^{-2}$ $7.41\times 10^{-3}$ $2.89\times 10^{-3}$ $2.98\times 10^{-3}$ $7.02\times 10^{-1}$
$f_{14}$ Star $2.34\times 10^{-1}$ $3.45\times 10^{-2}$ $2.34\times 10^{-1}$ $3.45\times 10^{-2}$ $1.83\times 10^{-1}$ $2.69\times 10^{-2}$ $8.82\times 10^{-1}$
Ring $1.68\times 10^{-1}$ $2.37\times 10^{-2}$ $1.68\times 10^{-1}$ $2.37\times 10^{-2}$ $1.54\times 10^{-1}$ $2.43\times 10^{-2}$ $8.90\times 10^{-1}$
Von Neumann $2.09\times 10^{-1}$ $2.38\times 10^{-2}$ $09\times 10^{-1} 2.2$$26\times 10^{-2}$ $1.72\times 10^{-1}$ $2.38\times 10^{-2}$ $9.12\times 10^{-1}$
$f_{15}$ Star $5.09\times 10^{-2}$ $9.97\times 10^{-3}$ $5.09\times 10^{-2}$ $9.97\times 10^{-3}$ $3.46\times 10^{-2}$ $1.05\times 10^{-2}$ $7.61\times 10^{-1}$
Ring $3.37\times 10^{-2}$ $1.04\times 10^{-2}$ $3.37\times 10^{-2}$ $1.04\times 10^{-2}$ $2.38\times 10^{-2}$ $9.25\times 10^{-3}$ $7.29\times 10^{-1}$
Von Neumann $4.42\times 10^{-2}$ $6.56\times 10^{-3}$ $4.32\times 10^{-2}$ $8.88\times 10^{-3}$ $3.08\times 10^{-2}$ $9.55\times 10^{-3}$ $7.76\times 10^{-1}$
$f_{16}$ Star $1.12\times 10^{-1}$ $42\times 10^{-2} 1.1$$12\times 10^{-1}$ $1.42\times 10^{-2}$ $1.02\times 10^{-1}$ $1.40\times 10^{-2}$ $6.85\times 10^{-1}$
Ring $1.09\times 10^{-1}$ $1.35\times 10^{-2}$ $1.08\times 10^{-1}$ $1.39\times 10^{-2}$ $1.02\times 10^{-1}$ $1.49\times 10^{-2}$ $7.56\times 10^{-1}$
Von Neumann $14\times 10^{-1} 1.1$$3\times 10^{-2} 1.1$$10\times 10^{-1}$ $1.26\times 10^{-2}$ $1.04\times 10^{-1}$ $1.47\times 10^{-2}$ $7.46\times 10^{-1}$
$f_{17}$ Star $1.78\times 10^{-1}$ $1.99\times 10^{-2}$ $1.78\times 10^{-1}$ $1.99\times 10^{-2}$ $1.53\times 10^{-1}$ $1.58\times 10^{-2}$ $8.45\times 10^{-1}$
Ring $1.51\times 10^{-1}$ $1.43\times 10^{-2}$ $1.49\times 10^{-1}$ $1.42\times 10^{-2}$ $37\times 10^{-1} 1.1$$15\times 10^{-2}$ $8.83\times 10^{-1}$
Von Neumann $1.68\times 10^{-1}$ $1.78\times 10^{-2}$ $1.67\times 10^{-1}$ $1.78\times 10^{-2}$ $1.47\times 10^{-1}$ $1.37\times 10^{-2}$ $8.54\times 10^{-1}$
$f_{18}$ Star $1.22\times 10^{-1}$ $2.04\times 10^{-2}$ $1.22\times 10^{-1}$ $2.04\times 10^{-2}$ $9.68\times 10^{-2}$ $1.54\times 10^{-2}$ $8.43\times 10^{-1}$
Ring $9.42\times 10^{-2}$ $1.69\times 10^{-2}$ $9.04\times 10^{-2}$ $1.53\times 10^{-2}$ $8.29\times 10^{-2}$ $1.39\times 10^{-2}$ $8.62\times 10^{-1}$
Von Neumann $1.16\times 10^{-1}$ $84\times 10^{-2} 1.1$$14\times 10^{-1}$ $1.81\times 10^{-2}$ $9.18\times 10^{-2}$ $1.47\times 10^{-2}$ $8.62\times 10^{-1}$
$f_{19}$ Star $1.60\times 10^{-1}$ $1.42\times 10^{-2}$ $1.60\times 10^{-1}$ $1.42\times 10^{-2}$ $1.41\times 10^{-1}$ $1.33\times 10^{-2}$ $6.57\times 10^{-1}$
Ring $1.41\times 10^{-1}$ $1.90\times 10^{-2}$ $1.41\times 10^{-1}$ $1.07\times 10^{-2}$ $1.29\times 10^{-1}$ $1.00\times 10^{-2}$ $5.72\times 10^{-1}$
Von Neumann $1.60\times 10^{-1}$ $1.65\times 10^{-2}$ $1.56\times 10^{-1}$ $1.33\times 10^{-2}$ $7\times 10^{-1} 1.33×1$$3\times 10^{-2}$ $6.77\times 10^{-1}$
$f_{20}$ Star $1.41\times 10^{-1}$ $1.20\times 10^{-2}$ $1.41\times 10^{-1}$ $1.20\times 10^{-2}$ $1.06\times 10^{-1}$ $3.12\times 10^{-2}$ $9.38\times 10^{-1}$
Ring $1.03\times 10^{-1}$ $2.57\times 10^{-2}$ $1.03\times 10^{-1}$ $2.57\times 10^{-2}$ $5.93\times 10^{-2}$ $5.09\times 10^{-2}$ $8.91\times 10^{-1}$
Von Neumann $1.35\times 10^{-1}$ $1.36\times 10^{-2}$ $1.34\times 10^{-1}$ $1.26\times 10^{-2}$ $9.53\times 10^{-2}$ $3.85\times 10^{-2}$ $9.28\times 10^{-1}$
$f_{21}$ Star $1.41\times 10^{-1}$ $4.53\times 10^{-2}$ $1.33\times 10^{-1}$ $1.99\times 10^{-2}$ $1.25\times 10^{-1}$ $2.60\times 10^{-2}$ $8.25\times 10^{-1}$
Ring $1.30\times 10^{-1}$ $2.59\times 10^{-2}$ $1.30\times 10^{-1}$ $2.59\times 10^{-2}$ $1.16\times 10^{-1}$ $2.03\times 10^{-2}$ $7.84\times 10^{-1}$
Von Neumann $1.55\times 10^{-1}$ $4.39\times 10^{-2}$ $1.51\times 10^{-1}$ $4.30\times 10^{-2}$ $1.27\times 10^{-1}$ $2.76\times 10^{-2}$ $8.25\times 10^{-1}$
$f_{22}$ Star $1.63\times 10^{-1}$ $7.86\times 10^{-2}$ $1.63\times 10^{-1}$ $6.67\times 10^{-2}$ $1.17\times 10^{-1}$ $3.93\times 10^{-2}$ $8.09\times 10^{-1}$
Ring $1.21\times 10^{-1}$ $4.88\times 10^{-2}$ $1.16\times 10^{-1}$ $3.15\times 10^{-2}$ $9.08\times 10^{-2}$ $2.64\times 10^{-2}$ $7.84\times 10^{-1}$
Von Neumann $1.38\times 10^{-1}$ $6.04\times 10^{-2}$ $1.38\times 10^{-1}$ $5.72\times 10^{-2}$ $1.08\times 10^{-1}$ $3.11\times 10^{-2}$ $8.23\times 10^{-1}$
$f_{23}$ Star $1.63\times 10^{-1}$ $5\times 10^{-2} 1.55×1$$5\times 10^{-1}$ $1.53\times 10^{-2}$ $1.57\times 10^{-1}$ $1.30\times 10^{-2}$ $7.19\times 10^{-1}$
Ring $1.63\times 10^{-1}$ $1.75\times 10^{-2}$ $61\times 10^{-1} 1.1$$16\times 10^{-2}$ $1.57\times 10^{-1}$ $1.24\times 10^{-2}$ $7.63\times 10^{-1}$
Von Neumann $1.62\times 10^{-1}$ $1.06\times 10^{-2}$ $53\times 10^{-1} 1.1$$13\times 10^{-2}$ $1.57\times 10^{-1}$ $1.25\times 10^{-2}$ $7.62\times 10^{-1}$
$f_{24}$ Star $5.38\times 10^{-2}$ $8.14\times 10^{-3}$ $5.31\times 10^{-2}$ $9.04\times 10^{-3}$ $4.41\times 10^{-2}$ $7.77\times 10^{-3}$ $6.58\times 10^{-1}$
Ring $4.29\times 10^{-2}$ $8.30\times 10^{-3}$ $4.02\times 10^{-2}$ $5.21\times 10^{-3}$ $3.68\times 10^{-2}$ $7.53\times 10^{-3}$ $7.57\times 10^{-1}$
Von Neumann $16\times 10^{-2} 5.5$$56\times 10^{-3}$ $5.02\times 10^{-2}$ $5.41\times 10^{-3}$ $4.28\times 10^{-2}$ $7.56\times 10^{-3}$ $7.49\times 10^{-1}$
6 Data-Driven and Interpretable Configuration Learning
In the IOHxplainer pipeline, extensive data is gathered on the behaviour and performance of various algorithm configurations for each topologies. While explainable AI (XAI) techniques can extract key insights from this data, there are further advantages to systematically storing and re-using it. In this section, we highlight several practical ways in which the collected data can be used to enhance the selection, configuration, and understanding of algorithm configurations.
Figure 8: Workflow for classifying PSO AOCC performance using an expanded set of ELA features
The IOHxplainer toolbox offers a valuable secondary benefit: it enables the straightforward training of machine learning models that can predict the most suitable algorithm configuration including both the choice of modules and their hyper-parameters based on observed characteristics of the optimization landscape. Figure 8 represents the workflow for performance classification using Exploratory Landscape Analysis (ELA) features. The diagram illustrates the process from feature extraction and augmentation to model training and evaluation for predicting PSO algorithm performance based on AOCC metrics. This is essentially the task of automated algorithm configuration and selection, echoing the goals of previous research such as those referenced in [24, 30] and [48]. What distinguishes our approach is that we merge the selection of the algorithm and the tuning of its configuration into a single, unified machine learning problem. We achieve this by employing multi-output models that can handle both classification (for discrete choices like module selection) and regression (for continuous settings like hyper-parameters) simultaneously. Specifically, our models are restricted to shallow Decision Tree (DT) [46] and Random Forest (RF) [39], chosen for their simplicity and interpretability. We gathered a wide range of Exploratory Landscape Analysis (ELA) features to show what we can do. To create these features, we set up an experiment where we sampled 1000 points from each instance in the BBOB suite. For each instance, we aimed for our machine learning models to find the best-performing algorithm configuration from our experimental results. Using models like RF and easy-to-understand DT, we can train our systems to predict which algorithm configurations will perform well based on the ELA features. This approach not only automates the process of algorithm selection and configuration, but also provides transparent and understandable decision rules thanks to the interpretable nature of the models used.
Figure 9: AOCCs of RF, DT, SB, and AB configurations for PSO using all topologies on all functions using LoFo. Models are trained excluding dimension 5.
Figure 10: AOCCs of RF, DT, SB, and AB configurations for PSO using all topologies on all functions using LoIo validation. Models are trained excluding dimension 5.
The AOCC performance of the Random Forest (RF) [39], Decision Tree (DT) [46], Single Best (SB), and Average Best (AB) configurations across all topologies and functions is presented in Figures 9 and 10. The key distinction is the validation strategy: Figure 9 tests generalization to unseen functions, and Figure 10 tests generalization to unseen problem instances. In Figures 11- 13, we observe how different PSO setup strategies work with Star, Ring, and Von Neumann patterns when applied to BBOB functions with d=5d{=}5, using both leave one function out (LoFo) and leave one instance out (LoIo) methods for validation. When looking at the Star topology, it’s clear that the RF model does much better than the others, achieving the lowest AOCC, whereas AB falls behind, showing the advantages of model-guided configurations. On the other hand, in the Ring topology, all three methods, AB, DT, and RF, perform similarly, with AB being a bit more stable and even holding its own against the model-based approaches. In the Von Neumann setup, AB consistently achieves the lowest AOCC, outperforming both DT and RF, which tend to be more variable. When it comes to validation methods, using LoIo often results in tighter and steadier AOCC distributions for every topology. This indicates that it does a better job of predicting how the model will perform with new instances. On the other hand, LoFo presents a tougher challenge for generalization and shows clearer differences in performance among the various methods. In general, model-based approaches seem to work best in Star topology, while AB tends to perform just as well or even better in more organized topologies like Ring and Von Neumann.
Figure 11: AOCC performance loss of PSO using Star topology on BBOB functions (d = 5), comparing RF, shallow DT, and AB configurations against the best single run, using LoFo (left) and LoIo (right) validation.
Figure 12: AOCC performance loss of PSO using Ring topology on BBOB functions (d = 5), comparing RF, shallow DT, and AB configurations against the best single run, using LoFo (left) and LoIo (right) validation.
Figure 13: AOCC performance loss of PSO using Von Neumann topology on BBOB functions (d = 5), comparing RF, shallow DT, and AB configurations against the best single run, using LoFo (left) and LoIo (right) validation.
6.1 Explaining PSO via Decision Trees and ELA Features
The decision trees for the Ring, Star, and Von Neumann topologies show how different landscape features impact the PSO algorithm’s performance at various inertia weights ($w=0.5,0.7,0.9$). For the Ring topology (Figure 14), the top feature is nbc.nb_fitness.cor, with ela_meta.quad_simple.cond coming in next. These features help categorize how the PSO performs based on different $w$ values. In the Star topology (Figure 15), nbc.nb_fitness.cor is also the main feature, but this time, features like disp.diff_mean_10 and ela.distr.skewness influence the decision-making process. As for the Von Neumann topology (Figure 16), the main feature here is disp.diff.mean_02, which is key for predictions, while ela_meta.lin_simple.coef.max_by_m and nbc.nn_nb.mean_ratio lend support. Each decision tree visually illustrates how instances are divided, and the leaf nodes show class distributions, pointing out the best weight settings for each topology.
Figure 14: Decision tree (depth=7) for inertia weights ($w$) module in PSO using Star Topology ($d=5$). The nodes show feature splits with value distributions- yellow: false, green: true. The black arrow marks the split threshold.
Figure 15: Decision tree (depth=7) for inertia weights ($w$) module in PSO using Ring Topology (d=5). Nodes show feature splits with value distributions—yellow: false, green: true. Black arrow marks split threshold.
Figure 16: Decision tree (depth=7) for inertia weights ($w$) module in PSO using Von Neumann Topology (d=5). Nodes show feature splits with value distributions—yellow: false, green: true. Black arrow marks split threshold.
Looking at the decision tree graphs, the Von Neumann layout stands out for its even and informative approach. It has important features like disp.diff.mean_02 that help sort PSO performance based on different inertia weights, hinting at its ability to adapt well and choose parameters more precisely. This adaptability allows Von Neumann to adjust smoothly to various challenges by recognizing small differences in features, helping the algorithm find the best settings for different situations. The Star layout, on the other hand, has a moderate level of complexity. It uses features such as nbc.nb_fitness.cor and ela.distr.skewness, which lead to quicker results but with less even class distributions, showing a bit of a trade-off between speed and strength. In contrast, the Ring layout is more straightforward, mainly relying on nbc.nb_fitness.cor, which means it has fewer splits and displays more uneven leaves. This setup results in a steadier but somewhat less adaptable performance. In general, Von Neumann seems to work well in different environments because it adapts easily. Star, on the other hand, tends to lead to quicker results but may lack some diversity. Ring provides a stable option, though it doesn’t adapt as well.
7 Conclusion
This work makes three key contributions to explainablity of PSO algorithm. First, we demonstrate the critical importance of landscape analysis through Exploratory Landscape Analysis (ELA), which provides fundamental insights into problem structure, multimodality, and difficulty characteristics. Our work shows how ELA features can guide algorithm selection and configuration by revealing intrinsic properties of optimization landscapes that influence heuristic performance. The IOHxplainer framework significantly enhances this analysis by enabling systematic investigation of hyperparameter contributions, particularly in quantifying how different PSO topologies interact with specific landscape features. Second, we apply this explainable approach to conduct the most comprehensive analysis of PSO topologies such as Star, Ring and Von Neumann configurations. The redesigned IOHxplainer toolbox turns out to be an invaluable asset in identifying how topological structure affects swarm behaviour, with Von Neumann consistently demonstrating superior time efficiency while Ring topology better maintains diversity in complex landscapes. Finally, we introduce a novel machine learning application by training Random Forest and Decision Tree classifiers on AOCC data. These models successfully predict optimal topology selection based on landscape characteristics, achieving accuracy. While focused on PSO, our methodology establishes a template for explainable analysis of any iterative swarm based meta-heuristic. Future directions include extending the framework to higher dimensions, integrating additional XAI techniques, and developing adaptive topology selection systems. This work advances optimization research from empirical benchmarking toward truly explainable algorithm design, with implications for both theoretical understanding and real-world applications.
Data availability
All research materials, including source code, experimental data, and detailed results, are publicly available to ensure transparency and reproducibility of our study. The complete implementation of the PSO algorithm with different topological configurations (Star, Ring, and Von Neumann) along with the benchmark functions and evaluation scripts can be accessed at: https://github.com/GitNitin02/ioh_pso
CRediT authorship contribution statement
Nitin Gupta:
Writing – original draft, Methodology, Experimentation. Bapi Dutta: Review & editing. Anupam Yadav: Writing – review & editing, Supervision.
Declaration of competing interest
The authors state that there are no competing interests to declare that might have influenced this research.
Acknowledgments
This study was supported by Dr. B. R. Ambedkar National Institute of Technology Jalandhar,